### v1

# Conformable Scaling and Critical Dynamics: A Unified Framework for Phase Transitions

Abstract
We investigate the application of conformable derivatives to model critical phenomena near continuous phase transitions. By incorporating a deformation parameter into the differential structure, we derive unified expressions for thermodynamic observables such as heat capacity, magnetization, susceptibility, and coherence length, each exhibiting power-law behavior near the critical temperature. The conformable derivative framework naturally embeds scale invariance and critical slowing down into the dynamics without resorting to fully nonlocal fractional calculus. Modified Ginzburg–Landau equations are constructed to model superconducting transitions, leading to analytical expressions for the order parameter and London penetration depth. Experimental data from niobium confirm the model’s applicability, showing excellent fits and capturing asymmetric scaling behavior around $T_{c}$. This work offers a bridge between classical mean-field theory and generalized scaling frameworks, with implications for both theoretical modeling and experimental analysis.
I Introduction
Critical phenomena play a central role in the understanding of phase transitions, where physical observables such as heat capacity, magnetization, and susceptibility exhibit non-analytic behavior characterized by power-law divergences near a critical temperature $T_{c}$. These behaviors are traditionally described through scaling hypotheses and renormalization group (RG) analysis, which explain the emergence of universality across diverse systems.
Phase transitions and critical phenomena represent one of the most fascinating and universal aspects of condensed matter physics, appearing across diverse systems from magnetic materials and superconductors to liquid-gas transitions and even biological systems Goldenfeld (1992); Stanley (1971). Near critical points, physical systems exhibit remarkable scaling behavior characterized by power-law divergences of thermodynamic quantities, with critical exponents that depend only on fundamental symmetries and spatial dimensionality rather than microscopic details Wilson (1971); Fisher and Barber (1972). This universality principle, established through decades of theoretical and experimental work, forms the cornerstone of modern critical phenomena theory.
The mathematical description of critical behavior has evolved significantly since the early phenomenological approaches of Landau and Ginzburg Landau (1965); Ginzburg and Landau (1965). The subsequent development of Ginzburg-Landau theory, particularly as systematized by de Gennes de Gennes (1966), provided a robust framework for understanding superconducting phase transitions and scaling behavior near critical points. The development of renormalization group theory by Wilson and others Wilson (1971); Wilson and Kogut (1974) provided deep insights into the origin of universality and scaling laws, while field-theoretic methods enabled precise calculations of critical exponents Zinn-Justin (2021). These advances established that critical phenomena emerge from the competition between thermal fluctuations and ordering tendencies, with correlation lengths diverging as $\xi\sim|T-T_{c}|^{-\nu}$ and other observables following characteristic power laws.
Despite these theoretical successes, many experimental systems exhibit deviations from ideal critical behavior due to finite-size effects, quenched disorder, non-equilibrium conditions, and other realistic complications Privman (1990); Harris (1974); Hohenberg and Halperin (1977). There remains a need for alternative frameworks that provide analytical tractability, flexible modeling, and connections to generalized thermodynamic formalisms. This is particularly relevant in systems where microscopic details are poorly known, long-range interactions are present, or nonlocal and memory effects become significant. Real materials often deviate from ideal critical behavior due to quenched disorder, finite-size effects, and non-equilibrium conditions. These deviations often manifest as modified scaling laws, anomalous relaxation dynamics, and effective critical exponents that differ from theoretical predictions. Understanding and modeling such complex critical behavior remains an active area of research with important implications for materials science, statistical physics, and complex systems.
In this work, we explore a deformation-based approach to critical phenomena using conformable derivatives Weberszpil and Helayël-Neto (2015, 2016); Weberszpil and Sotolongo-Costa (2017); Weberszpil and Chen (2017); Sotolongo-Costa and Weberszpil (2021); Rosa and Weberszpil (2018); Godinho et al. (2020), a form of generalized calculus that introduces a deformation parameter $\mu$ into the derivative operator:
$D_{T}^{(\mu)}f(T):=T^{1-\mu}\frac{df}{dT}.$ (1)
This operator naturally yields temperature-dependent scaling behavior and introduces power-law features without requiring full fractional integration or nonlocal convolution terms. It thus offers an intermediate framework between classical and fractional derivatives, retaining analytical simplicity while capturing essential scaling features.
We apply this formalism to describe thermodynamic observables near continuous phase transitions, including heat capacity, magnetization, susceptibility, and correlation length. In each case, we derive critical exponents as functions of the deformation parameter and temperature, leading to expressions compatible with known mean-field results. Furthermore, we extend the method to superconducting phase transitions by constructing a modified Ginzburg–Landau (GL) equation with conformable kinetic terms. This allows us to analyze the temperature dependence of the superconducting order parameter, London penetration depth, and heat capacity within the same mathematical framework.
With this approach we claim that it is possible to obtain a unified formalism for deriving all major critical exponents; analytical expressions that match experimental data with high fidelity; a bridge between classical mean-field theory and generalized (e.g., Tsallis) statistical mechanics and also a minimal extension of known differential equations without resorting to nonlocal fractional operators.
By maintaining dimensional consistency and compatibility with equilibrium thermodynamics, the conformable derivative framework provides a useful and versatile tool for exploring critical dynamics in both classical and quantum systems. The results presented here demonstrate the capacity of this formalism to reproduce known scaling laws while offering new insights into generalized scaling behavior near criticality.
We derive unified expressions for critical exponents in terms of conformable parameters that emerge from these physical mechanisms, demonstrating that the approach extends rather than contradicts renormalization group theory. The formalism is applied to magnetic phase transitions, superconducting transitions, and other critical phenomena, with specific predictions for experimental verification.
II Conformable Derivatives in Physical Systems
Near critical points, physical systems exhibit critical slowing down, where the relaxation time $\tau$ diverges as the correlation length $\xi$ increases. This behavior is characterized by the dynamic scaling law
$\tau\sim\xi^{z}\sim|T_{c}-T|^{-z\nu},$ (2)
where $z$ is the dynamic critical exponent and $\nu$ is the correlation length exponent Hohenberg and Halperin (1977). $\nu$ is the correlation length exponent from $\xi(T)\sim|T-T_{c}|^{-\nu}$.
Observable thermodynamic quantities exhibit characteristic power-law behavior in the vicinity of continuous (second-order) phase transitions. Specifically, the heat capacity diverges as
$C(T)\sim|T_{c}-T|^{-\alpha},$
where $\alpha$ is the critical exponent associated with the specific heat. The spontaneous magnetization follows
$|M(T)|\sim(T_{c}-T)^{\beta},\quad\text{for }T<T_{c},$
with $\beta$ denoting the critical exponent for the order parameter. The magnetic susceptibility diverges according to
$\chi(T)\sim|T_{c}-T|^{-\gamma},$
where $\gamma$ is the critical exponent characterizing the divergence of the susceptibility. These exponents $\alpha$, $\beta$, and $\gamma$ define universal scaling behavior near the critical point and are largely independent of the microscopic details of the system. We will show that it naturally introduces temperature-weighted derivatives that match the scaling behavior of critical observables. This property embeds scale invariance directly into the differential structure of the equations, enabling straightforward generation of power-law solutions.
The kinetic coefficient $\Gamma(T)$ plays a central role in the time-dependent relaxation dynamics of systems near criticality. It essentially measures how rapidly a system returns to equilibrium after a small perturbation. Physically, it serves as a mobility-like parameter, often appearing in the Langevin equation or time-dependent Ginzburg–Landau models as a prefactor to the functional derivative of the free energy. Since the kinetic coefficient $\Gamma(T)$ is inversely proportional to the relaxation time, it inherits the corresponding divergence:
$\Gamma(T)\sim|T_{c}-T|^{z\nu}.$ (3)
This reflects the critical slowing down of dynamics near $T_{c}.$
This decreases as $T\rightarrow Tc,$reflecting the critical slowing down of dynamics: the relaxation time diverges, and the kinetic response $\Gamma(T)$ vanishes. In over-damped systems, this corresponds to a loss of mobility, where the system becomes increasingly inert to perturbations near the critical point.
This work addresses the fundamental question of how conformable derivatives emerge naturally in critical phenomena while offering genuine physical insights beyond classical mean-field theory. We establish rigorous connections between the conformable derivative formalism and established physical mechanisms. First, we demonstrate how anomalous relaxation dynamics near the critical temperature $Tc,$give rise to conformable structures in the adiabatic limit. Second, we show that finite-size effects and boundary constraints lead to temperature-dependent prefactors consistent with the conformable framework. Third, we connect the presence of quenched disorder and spatial inhomogeneities to the emergence of modified scaling laws captured effectively by conformable derivatives. Finally, we relate the conformable exponent to underlying fractal geometry, interpreting it as an effective dimension governing critical fluctuations.
Remark on the use of critical scaling forms:
When expressing power-law behavior near the critical temperature $T_{c}$, the choice of expression depends on the temperature regime. For temperatures below the critical point ($T<T_{c}$), one uses $(T_{c}-T)^{\beta}$, as this quantity remains positive and correctly describes observables such as the order parameter (e.g., magnetization or superconducting gap), which typically grow as $T$ decreases below $T_{c}$. For temperatures above the critical point ($T>T_{c}$), the appropriate form is $(T-T_{c})^{\beta}$, which applies to quantities that diverge as the system approaches $T_{c}$ from above, such as the heat capacity or magnetic susceptibility. When a symmetric form is desired, valid on both sides of the transition, a common convention is to use the absolute value $|T-T_{c}|^{\beta}$, which appears frequently in generalized scaling laws and in the unified description of critical divergences.
Conformable Derivatives: Bridging Classical, Geometric, and Fractional Perspectives
The conformable derivative provides a powerful intermediary between classical differential operators and fully nonlocal fractional derivatives. While fractional calculus introduces long-range memory and non-locality, it is often analytically cumbersome and dimensionally ambiguous. The conformable derivative, as defined in Eq.(1), offers a tractable alternative. For $\mu=1$, it reduces to the classical derivative, while for $\mu<1$, it captures essential features of fractional scaling, such as subdiffusion, intermittency, and fractal geometry, without requiring nonlocal integration kernels.
This makes the conformable framework especially suitable for systems exhibiting complex thermodynamic behaviors, including those described by generalized statistics (e.g., Tsallis entropy). Crucially, from a dimensional standpoint, the conformable derivative maintains physical unit consistency:
$[D_{T}^{(\mu)}f]=\frac{[f]}{[T]^{\mu}},$ (4)
and all critical exponents derived from it (e.g., $\alpha,\beta$) remain dimensionless. Modified Ginzburg-Landau equations incorporating conformable terms remain compatible with thermodynamic scaling laws and unit balance, confirming both mathematical consistency and physical viability.
Beyond this formal consistency, the conformable derivative admits a compelling geometric interpretation. The operator $D_{T}^{(\mu)}$ effectively deforms the thermal geometry by introducing a temperature-weighted directional response:
$\text{Effective response}\propto T^{1-\mu}\frac{d}{dT}.$ (5)
This can be viewed as a curvature in the space of thermal states, analogous to how spacetime curvature in general relativity modifies geodesics. Here, the thermal evolution is bent by the conformable weighting, making the system’s response scale-sensitive. This captures phenomena like anomalous diffusion, divergence of correlation length, or emergent power-law scaling near criticality.
Importantly, the framework retains analytical tractability. Unlike renormalization group treatments, conformable models yield closed-form or semi-analytic expressions for key observables such as the order parameter $\psi(T)$, London penetration depth $\lambda_{L}(T)$, and heat capacity $C(T)$. These expressions are not only dimensionally coherent but also directly fit experimental data. In particular, fits to superconducting phase transitions recover expected mean-field exponents such as $\beta\approx 1/2$, thereby validating the model’s empirical utility.
In summary, the conformable derivative offers a physically motivated, analytically manageable, and geometrically meaningful framework for modeling critical phenomena. It seamlessly bridges classical dynamics, geometric deformation, and fractional scaling, making it a promising tool for exploring universality, non-locality, and thermodynamic complexity in condensed matter systems and beyond.
To explore how deformed calculus affects thermodynamic behavior near criticality, we now introduce the conformable derivative framework and show how it modifies standard scaling relations.
III Critical Exponents and Thermodynamic Quantities
In this section, we derive the scaling behavior of key thermodynamic quantities near the critical temperature $T_{c}$, using the conformable derivative framework. This approach yields expressions for the critical exponents $\alpha$, $\beta$, $\gamma$, and $\nu$, which characterize the singular behavior of heat capacity, magnetization, magnetic susceptibility, and correlation length, respectively.
We begin by applying this operator to each observable and derive the corresponding critical exponent in terms of the deformation parameter $\mu$ and characteristic coefficients. These expressions will form the foundation for the quantitative modeling of critical behavior in later sections.
III.1 Phenomenological Origin of the Conformable Critical Equation - Heat Capacity
To model power-law divergences near critical points, it is common to describe thermodynamic observables such as susceptibility or concentration fluctuations with first-order differential equations whose solutions yield singularities at the critical temperature $T_{c}$. In the classical framework, such behavior can be captured by equations of the form:
$\frac{dC}{dT}=-\frac{\alpha}{T_{c}-T}C(T),$ (6)
which yields the well-known scaling law:
$C(T)\sim|T_{c}-T|^{-\alpha}.$ (7)
To incorporate generalized dynamics, such as memory effects, anomalous relaxation, or scaling violations associated with nonextensive systems, we replace the classical derivative with the conformable derivative of order $\mu_{C}\in(0,1]$, defined as:
$D_{T}^{(\mu_{C})}f(T):=T^{1-\mu_{C}}\frac{df}{dT}.$ (8)
This local deformation introduces intrinsic scaling into the dynamics while preserving analytic solvability.
We thus propose the following phenomenological equation:
$D_{T}^{(\mu_{C})}C(T)=-\kappa(T_{c}-T)^{-1}C(T),$ (9)
where $\kappa$ is a dimensionless constant governing the strength of the singularity.
Substituting the definition of the conformable derivative, the equation becomes:
$T^{1-\mu_{C}}\frac{dC}{dT}=-\kappa(T_{c}-T)^{-1}C(T).$ (10)
Dividing both sides by $C(T)$ and integrating:
[…]
Near the critical point $T\approx T_{c}$, we approximate $T^{\mu_{C}-1}\approx T_{c}^{\mu_{C}-1}$, obtaining:
$\ln C(T)\sim\kappa T_{c}^{\mu_{C}-1}\ln|T_{c}-T|,$ (12)
and hence the critical behavior:
$C(T)\sim|T_{c}-T|^{-\alpha},\quad\text{with}\quad\alpha=\kappa T_{c}^{\mu_{C}-1}.$ (13)
This formulation provides a conformable generalization of classical critical scaling laws, embedding the power-law divergence directly into the modified dynamics. It is particularly suitable for systems exhibiting anomalous thermodynamic behavior, nonlocality, or deviations from classical universality classes.
This yields the explicit critical form of the heat capacity near the second-order phase transition:
$C_{V}(T)=B\,|T_{c}-T|^{-\alpha},$ (14)
where $B$ is a positive amplitude constant.
The empirical data from physical systems such as niobium often exhibit asymmetric critical behavior, where the divergence of $C_{V}$ is different below and above the critical temperature. This asymmetry arises from distinct microscopic mechanisms in the ordered (superconducting) and disordered (normal) phases.
To account for this, we adopt a piecewise power-law model for the heat capacity:
$C_{V}(T)=\begin{cases}B_{1}(T_{c}-T)^{-\alpha_{1}},&T<T_{c}\\ B_{2}(T_{c}-T)^{-\alpha_{2}},&T>T_{c}\end{cases}$ (15)
where $B_{1}$, $B_{2}$, $\alpha_{1}$, and $\alpha_{2}$ are fitting parameters.
In more specific details, the piecewise structure of Eq. (15) reflects the physically distinct regimes on either side of the critical temperature $T_{c}$. In many experimental systems, including niobium, the divergence of the heat capacity is not symmetric with respect to $T_{c}$. That is, the scaling exponent and amplitude below the transition ($T<T_{c}$) differ from those above it ($T>T_{c}$). This asymmetry is rooted in microscopic differences between the ordered (superconducting) and disordered (normal) phases, and is supported by both theoretical models and experimental heat capacity curves Stanley (1971). Therefore, we adopt a piecewise formulation, which allows distinct critical exponents $\alpha_{1}$, $\alpha_{2}$ and prefactors $B_{1}$, $B_{2}$ to be fitted on either side of the transition. Here the parameters $\alpha_{i}=\kappa_{i}T_{c}^{\mu_{C,i}-1}$, and the prefactors $B_{1}$ and $B_{2}$ account for the asymmetry of the heat capacity below and above the transition.
With these physical justifications, we claim that the piecewise expression provides a better fit to real data than a symmetric model and remains analytically tractable within the conformable framework.
Also, for computational purposes, to avoid divergence exactly at $T=T_{c}$, we introduce a regularization parameter $\epsilon>0$, resulting in the smoothed model:
$C_{V}^{\text{reg}}(T)$ =$=\begin{cases}B_{1}\left(|T_{c}-T|+\epsilon\right)^{-\alpha_{1}},&T<T_{c}\\[6.0pt] B_{2}\left(|T_{c}-T|+\epsilon\right)^{-\alpha_{2}},&T>T_{c}\end{cases}$ (16)
In what follows, each Tc is relative to the phenomenon treated and corresponding. We keep the same symbol to avoid overloading the notation.
III.2 Magnetization
Now let’s analise the magnetization with application of the deformed operator.
It is well now that the magnetization behaves as:
$|M|\sim(T_{c}-T)^{\beta}.$ (17)
Now, considering the application of the conformable derivative, gives:
$D_{T}^{(\mu_{M})}|M|=\gamma(T_{c}-T)^{-1}|M|,$ (18)
where the deformed derivative is $D_{T}^{(\mu_{M})}f(T)=T^{1-\mu_{M}}\frac{df(T)}{dT}$.
The resulting differential expression can be written as:
$\frac{d|M|}{|M|}=\gamma T^{\mu_{M}-1}(T_{c}-T)^{-1}\,dT$ (19)
We integrate both sides. The left-hand side yields:
$\int\frac{d|M|}{|M|}=\ln|M|+\text{const.}$. (20)
For the right-hand side, we consider the behavior close to the critical temperature $T\to T_{c}^{-}$. Assuming that $T^{\mu_{M}-1}$ varies slowly compared to the divergence at $T=T_{c}$, we approximate:
$T^{\mu_{M}-1}\approx T_{c}^{\mu_{M}-1}$ (21)
Hence, the integral becomes:
[…]
$\approx$ $\gamma T_{c}^{\mu_{M}-1}\int\frac{dT}{T_{c}-T}=$ (23)
$=$ $-\gamma T_{c}^{\mu_{M}-1}\ln(T_{c}-T).$ (24)
Substituting back, we obtain:
$\ln|M|$ $\approx-\gamma T_{c}^{\mu_{M}-1}\ln(T_{c}-T),$ (25)
witch leads to
$|M|$ $\sim(T_{c}-T)^{\gamma T_{c}^{\mu_{M}-1}}.$ (26)
Identifying the standard critical scaling form $|M|\sim(T_{c}-T)^{\beta}$, we finnaly that the exponet $\beta$ is:
$\beta=\gamma T_{c}^{\mu_{M}-1}.$ (27)
This expression relates the critical exponent $\beta$ to the temperature-scaling exponent $\mu_{M}$ and the kinetic prefactor $\gamma$, providing a natural thermodynamic link between dynamic behavior and critical ordering.
III.3 Magnetic Susceptibility
Magnetic susceptibility diverges at criticality:
$\chi(T)\sim|T_{c}-T|^{-\gamma}$ (28)
This power-law has been treated in both classical and generalized frameworks Tsallis (1988); Plastino and Plastino (1993).
Following the same approach used for the calculation of heat capacity an magnetization, we can write the dedormed differential equation fo magnetic susceptibility as:
$D_{T}^{(\mu_{\chi})}\chi(T)=-\lambda(T_{c}-T)^{-1}\chi(T)$ (29)
Again, we can conclude that $\gamma$ exponent is given by
$\gamma=\lambda T_{c}^{\mu_{\chi}-1}.$ (30)
III.4 Correlation Length
The correlation length also diverges as
$\xi(T)\sim|T_{c}-T|^{-\nu}$ (31)
Following the approach, we can cast $\xi(T),$
$D_{T}^{(\mu_{\xi})}\xi(T)=-\rho(T_{c}-T)^{-1}\xi(T),$ (32)
leading to the relation for $\nu$:
$\nu=\rho T_{c}^{\mu_{\xi}-1}.$ (33)
This behavior is central to the renormalization group picture of criticality Goldenfeld (1992) and connects to the scaling approaches pioneered by de Gennes de Gennes (1979) in various critical systems.
IV Unified Framework and Theoretical Implications
Building on the conformable formalism, we derive modified evolution equations for thermodynamic observables and extract analytic expressions for the associated critical exponents.
The use of conformable derivatives to model critical phenomena offers theoretical and practical advantages:
IV.1 Unified Treatment of Critical Exponents
The model provides a single formalism for expressing all major critical exponents:
$\alpha=\kappa T_{cC}^{\mu_{C}-1},\quad\beta=\gamma T_{cM}^{\mu_{M}-1},\quad\gamma=\lambda T_{c\chi}^{\mu_{\chi}-1},\quad\nu=\rho T_{c\xi}^{\mu_{\xi}-1},$ (34)
where we introduced in this line a ddistinct notation for the critical temperature related to with each physical quantity.
The approach used here avoids the need for distinct phenomenological models for each observable, offering a unified perspective grounded in deformed calculus.
IV.2 Unified Thermodynamic Framework
The use of deformed derivatives provides a natural bridge between classical critical phenomena and generalized thermodynamics, particularly within the Tsallis statistical framework Tsallis (1988); Plastino and Plastino (1993). By incorporating temperature-dependent scaling, long-range correlations, memory, and fractal-like structures, the model effectively captures behaviors beyond the scope of classical universality classes.
A major advantage of this approach is its analytic tractability. In contrast to renormalization group techniques, which often require complex iterative procedures, the present model leads to closed-form expressions for critical behavior. The resulting differential equations are easily solvable and lend themselves to direct comparison with experimental data, making the method valuable for both theorists and experimentalists.
Moreover, the conformable or scale-deformed derivatives employed in the formulation naturally introduce power-law solutions. This mirrors the observed scaling of physical observables near the critical temperature $T_{c}$, where critical exponents emerge from the underlying structure of the modified dynamics rather than requiring external assumptions.
By using deformed derivatives, the model connects naturally to generalized thermodynamics, particularly Tsallis statistics Tsallis (1988); Plastino and Plastino (1993). It captures effects from long-range correlations, memory, and fractal-like structures, making it well-suited for systems that deviate from classical universality.
Unlike renormalization group methods, this model yields closed-form solutions and is analytically manageable. Its equations can be solved directly or fitted to data without computational overhead, making it practical for both theorists and experimentalists.
The conformable derivative itself introduces power-law behavior, which matches the scaling properties of observables near $T_{c}$.
To benchmark the conformable model, we now compare its predictions to classical mean-field theory and investigate its behavior in the limiting case where the deformation vanishes.
V Modified GL Equation
To derive the Ginzburg–Landau (GL) equation with a conformable derivative and nonlinear interaction term, we remember first the conventional GL theory. As treated by de Gennes de Gennes (1966), the free energy functional is:
$\mathcal{F}[\psi]=\int d^{d}r\left[a(T)|\psi(\mathbf{r})|^{2}+\frac{b}{2}|\psi(\mathbf{r})|^{4}+\frac{1}{2m}\left|\nabla\psi(\mathbf{r})\right|^{2}\right],$ (35)
The classical Ginzburg–Landau theory describes the thermodynamic behavior of a complex order parameter $\psi(\mathbf{r})$ near the critical temperature $T_{c}$. The GL free energy functional is given by: where:
  • $\psi(\mathbf{r})$: complex order parameter (e.g., superconducting wavefunction),
  • $a(T)=a_{0}(T_{c}-T)$: temperature-dependent coefficient (with $a_{0}>0$),
  • $b>0$: stabilizing coefficient ensuring boundedness of the free energy,
  • $m$: a generalized mass or stiffness parameter,
  • $\nabla\psi$: spatial gradient of the order parameter,
  • $d$: spatial dimensionality of the system.
The classical Ginzburg–Landau theory describes the thermodynamic behavior of a complex order parameter $\psi(\mathbf{r})$ near the critical temperature $T_{c}$. The GL free energy functional is given by the Eq. (35).
To determine the equilibrium configuration, we minimize $\mathcal{F}[\psi]$ with respect to $\psi^{*}(\mathbf{r})$, $\frac{\delta\mathcal{F}}{\delta\psi^{*}(\mathbf{r})}=0,$ leading to the Euler–Lagrange equation:
$a(T)\psi(\mathbf{r})+b|\psi(\mathbf{r})|^{2}\psi(\mathbf{r})-\frac{1}{2m}\nabla^{2}\psi(\mathbf{r})=0.$ (36)
This nonlinear partial differential equation governs the spatial behavior of the order parameter near criticality. In the case of a spatially uniform system (i.e., $\nabla\psi=0$), this reduces to the simpler algebraic equation:
$a(T)\psi+b|\psi|^{2}\psi=0,$ (37)
whose solutions describe the second-order phase transition and yield the critical exponent $\beta=\tfrac{1}{2}$ under mean-field conditions. This will be confirmed in the sequence.
In the classical Ginzburg–Landau (GL) theory, the behavior of the order parameter $\psi(T)$ near the critical temperature $T_{c}$ follows the scaling law:
$|\psi(T)|\sim(T_{c}-T)^{\beta},\qquad\text{for }T<T_{c},$ (38)
with the critical exponent $\beta=\tfrac{1}{2}$ under mean-field conditions. This result arises from minimizing the standard GL free energy functional.
To justify this result, we assume that the system is in thermodynamic equilibrium, the homogeneity and locality (no spatial gradients or long-range correlations), the fluctuations are negligible, the free energy is analytic near $T_{c}$ and that the order parameter $\psi$ is uniform and real.
In the full Ginzburg–Landau free energy functional, gradient term $\frac{1}{2m}\left|\nabla\psi(\mathbf{r})\right|^{2}$are included to account for the energy cost associated with variations of the order parameter $\psi$ in time, temperature, or space. This terms is essential in dynamic or spatially inhomogeneous contexts.
However, in the classical derivation of the critical exponent $\beta=\tfrac{1}{2}$, we work under a specific set of equilibrium assumptions that the system is in thermodynamic equilibrium, the order parameter $\psi$ is spatially and temporally uniform and there are no external fields or gradients applied.
We focus on minimizing the free energy to find static equilibrium configurations.
Under these conditions, the order parameter does not vary with position ( or even with temperature) in a local region, $\nabla\psi(\mathbf{r})=0.$
As a result, the gradient terms vanish, $\left|\nabla\psi(\mathbf{r})\right|^{2}=0,$ and under these conditions, the time-independent, spatially uniform GL free energy density becomes:
$\mathcal{F}[\psi]=a(T)|\psi|^{2}+\frac{b}{2}|\psi|^{4},$ (39)
which depends solely on the magnitude of $\psi$ and the temperature $T$.
Here
$a(T)=a_{0}(T_{c}-T),\qquad b>0.$ (40)
To minimize the free energy, we take the derivative with respect to $\psi^{*}$ (or $\psi$, since it’s real) and set it to zero:
$\frac{d\mathcal{F}}{d\psi}=2a(T)\psi+2b|\psi|^{2}\psi=0.$ (41)
This yields:
$\psi\left[a(T)+b|\psi|^{2}\right]=0.$ (42)
To gain insight about deformed options, we remember some anlisys. There are two solutions:
  1. 1.
    $\psi=0$, which corresponds to the disordered (symmetric) phase,
  2. 2.
    $|\psi|^{2}=-\frac{a(T)}{b}$, which corresponds to the ordered (broken-symmetry) phase for $a(T)<0$, i.e., $T<T_{c}$.
  3. 3.
    We now assume a form $a(T)=a_{0}(T_{c}-T)$.
Below $T_{c}$, we have $a(T)<0$, so the nontrivial solution becomes:
$|\psi|^{2}=-\frac{a_{0}(T_{c}-T)}{b}=\frac{a_{0}}{b}(T_{c}-T).$ (43)
Taking the square root, we otain:
$|\psi(T)|=\sqrt{\frac{a_{0}}{b}}\cdot(T_{c}-T)^{1/2}.$ (44)
Thus, the order parameter vanishes as:
$|\psi(T)|\sim(T_{c}-T)^{\beta},\quad\text{with }\beta=\frac{1}{2},$ (45)
$T<T_{C}.$
This result is the well known mean-field prediction for the order parameter exponent and is valid under the assumptions stated above. In lower dimensions ($d\leq 3$), fluctuations become important, and this mean-field result typically underestimates the true critical exponent.
VI The Conformable GL Model.
With the general framework established, we next apply it to specific thermodynamic quantities—namely, specific heat and magnetic susceptibility—highlighting the modifications induced by the conformable deformation.
In the standard Ginzburg–Landau (GL) theory, spatial variations of the order parameter $\psi(\mathbf{r})$ contribute to the free energy via the gradient term $\frac{1}{2m}\left|\nabla\psi(\mathbf{r})\right|^{2}$, which penalizes sharp spatial inhomogeneities which penalizes sharp spatial inhomogeneities and plays a key role in capturing coherence, textures, and defects. This is essential in systems where the spatial coherence or texture of the order parameter matters, such as vortices in superconductors.
In our modified GL framework, however, we are concerned with the thermal evolution of the order parameter $\psi(T)$, describe how the this order parameter evolves as a function of temperature $T$, especially in the vicinity of the critical point (critical temperature) $T_{c}$. That is, we study how the magnitude of $\psi$ changes as a function of temperature, under the assumption that the system is homogeneous in space.
To build a thermodynamic analog of the GL theory, we formally treat temperature $T$ as the independent coordinate, akin to space $\mathbf{r}$ in the conventional setting. In this case, we introduce a conformable (deformed) derivative with respect to temperature, denoted by:
$T_{\alpha}\psi(T):=T^{1-\alpha}\frac{d\psi}{dT},$ (46)
which naturally embeds temperature-dependent scaling into the dynamics. This operator acts as a thermal gradient and captures power-law behavior typical of critical phenomena.
This operator generalizes the usual derivative and introduces a natural temperature-dependent scaling, allowing the model to capture anomalous thermodynamic behavior while remaining analytically tractable. The corresponding modified GL free energy becomes
[…]
where $a(T)=a_{0}(T-T_{c})$ encodes the proximity to the critical point, $b>0$ ensures stability, and the last term mimics the kinetic contribution, now expressed in thermal rather than spatial form. The parameter $m$ retains its interpretation as a stiffness constant, and $\alpha\in(0,1]$ quantifies the deviation from classical scaling. For $\alpha=1$, the standard GL theory is recovered.
This formulation preserves the physical dimensions of energy density and ensures that all critical exponents derived from it are dimensionless and compatible with thermodynamic scaling theory. Furthermore, the temperature-dependent weighting introduced by $T^{1-\alpha}$ can be interpreted geometrically as a curvature in thermal space, analogous to how curved metrics affect geodesics in general relativity. It effectively embeds power-law sensitivity into the response functions of the system and models thermal evolution in a scale-aware manner. By replacing spatial derivatives with conformable thermal derivatives, the framework captures generalized thermodynamic behavior, such as fractality, memory, and scale-invariant dynamics, without invoking fully nonlocal fractional calculus, thereby offering both mathematical simplicity and physical relevance.
To minimize the free energy, we use the variational principle $\delta\mathcal{F}_{\alpha}/\delta\psi^{*}=0$, which yields the modified GL equation:
$a(T)\psi+b|\psi|^{2}\psi-\frac{1}{2m}T^{2(1-\alpha)}\frac{d^{2}\psi}{dT^{2}}-\frac{(1-\alpha)}{m}T^{1-2\alpha}\frac{d\psi}{dT}=0,$ (48)
where $\beta$ is a critical exponent depending on $\alpha$, and $\psi_{0}$ is the amplitude. This expression recovers the expected behavior from mean-field theory when $\alpha=1$, but allows deviations due to the deformation parameter.
These equations can also be obtained via a variational formulation adapted to deformed derivatives, as discussed in Weberszpil and Helayël-Neto (2016).
The nonlinear term $|\psi|^{2}\psi$ contributes to saturation of the order parameter as $T\rightarrow 0$, while the conformable kinetic term introduces corrections near $T_{c}$, mimicking effects from spatial or temporal fluctuations.
This formulation enables capturing intermediate critical dynamics and non-locality without invoking full fractional calculus or spatial disorder (Tinkham (2004); Bardeen et al. (1957); Gorter and Casimir (1934); Park and Dorsey (2003)), This formulation enables capturing intermediate critical dynamics and non-locality without invoking full fractional calculus or spatial disorder, making it suitable for describing second-order phase transitions such as superconductivity.
This approach complements traditional models such as the BCS theory Bardeen et al. (1957), the two-fluid model by Gorter and Casimir (1934), and the phenomenological framework detailed in Tinkham (2004). For systems with strong fluctuations and short coherence length, additional comparison with the work of Park and Dorsey Park and Dorsey (2003) provides insight into limitations and refinements. and non-locality without invoking full fractional calculus or spatial disorder, This formulation enables capturing intermediate critical dynamics and non-locality without invoking full fractional calculus or spatial disorder, making it suitable for describing second-order phase transitions such as superconductivity.
A key quantity in critical phenomena is the correlation length, whose divergence governs universality. We now derive its scaling law within the conformable approach and discuss physical implications.
VII Application to Superconducting Phase Transitions
Superconducting phase transitions, especially in type-II and high-$T_{c}$ materials, are characterized by critical phenomena that can be reformulated using conformable derivatives. This includes the temperature behavior of the order parameter, coherence length, penetration depth, and heat capacity.
VII.1 Simplified Dynamics Near the Critical Point: Connection to the Conformable Derivative Framework
We begin with the full conformable Ginzburg–Landau equation obtained by minimizing the deformed free energy functional, to obtain the 48.
The nonlinear second-order differential equation governs the thermal evolution of the order parameter $\psi(T)$, incorporating both nonlinear saturation ($|\psi|^{2}\psi$) and conformable kinetic effects through temperature-dependent coefficients.
Near the critical point, however, the system undergoes a continuous phase transition. In this regime, we make the following approximations to simplify the analysis:
  1. 1.
    Assume $\psi(T)$ is real and slowly varying. This allows us to treat $\psi$ and its derivatives in a simplified manner.
  2. 2.
    Neglect higher-order kinetic corrections. The second-order derivative term and the prefactor corrections from conformable calculus are of higher order and can be dropped close to $T_{c}$. This reduces the second-order equation to a first-order approximation.
  3. 3.
    Retain only the dominant balance between the kinetic term and the potential terms.
Taking into acout the above approximations, the reduced time-dependent equation can be cast as
$\frac{d\psi}{dT}=a(T)\psi-b|\psi|^{2}\psi,$ (49)
where $a(T)=a_{0}(T_{c}-T)$, and $b>0$ ensures stability of the ordered phase.
VII.2 Emergence of Critical Scaling from Modified GL Relaxation with Thermally Deformed Kinetics
While the standard Ginzburg-Landau (GL) equation near the critical temperature $T_{c}$ yields regular (non-singular) behavior for the order parameter $\psi(T)$, a phenomenological equation capable of reproducing power-law critical behavior can be obtained by introducing a thermally deformed kinetic coefficient into the GL-like relaxation framework. This modification effectively incorporates the temperature-dependent slowing down of dynamics near $T_{c}$ into the evolution of the order parameter.
As before, we generalize the time derivative to a conformable derivative:
$\frac{d\psi}{dT}\longrightarrow D_{T}^{(\mu_{\psi})}\psi(T):=T^{1-\mu_{\psi}}\frac{d\psi}{dT},$ (50)
which reflects the modified temperature response of the system under conformable thermodynamic dynamics.
At the heart of the conformable dynamics approach lies a generalized time-dependent Ginzburg–Landau-type equation that governs the thermal evolution of the order parameter $\psi(T)$ under non-equilibrium conditions. This equation introduces a thermally modulated kinetic response and is given by
$\Gamma(T)\,T^{1-\mu_{\psi}}\frac{d\psi}{dT}$ $=-\frac{\delta\mathcal{F}}{\delta\psi},$ (51)
where $\Gamma(T)$ is a temperature-dependent kinetic coefficient encoding critical slowing down, $\mu_{\psi}$ is the conformable deformation parameter, and $\mathcal{F}[\psi]$ is the free energy functional. This relation is not merely a phenomenological extension but constitutes the cornerstone of the proposed non-equilibrium conformable dynamics framework. It encapsulates the scale-sensitive kinetic behavior characteristic of critical systems, embedding power-law relaxation directly into the governing evolution law. Equation (51) allows the derivation of critical exponents such as $\beta$ from first principles within a thermodynamically consistent and geometrically interpretable formalism, thereby bridging equilibrium mean-field results with more general, deformation-induced scaling behaviors.
We take the standard GL potential (neglecting gradient terms):
[…]
which gives the variational derivative:
$\frac{\delta\mathcal{F}}{\delta\psi}=a_{0}(T_{c}-T)\psi+b|\psi|^{2}\psi.$ (53)
Substituting into the relaxation equation 51 yields:
$\Gamma(T)T^{1-\mu_{\psi}}\frac{d\psi}{dT}=-a_{0}(T_{c}-T)\psi-b|\psi|^{2}\psi.$ (54)
Near the critical point $T\to T_{c}$, the order parameter $\psi(T)$ becomes small, allowing us to neglect the nonlinear term in Eq. (54). The relaxation equation then reduces to a linearized form:
$\Gamma(T)\,T^{1-\mu_{\psi}}\frac{d\psi}{dT}\approx-a_{0}(T_{c}-T)\psi.$ (55)
Rewriting this in differential form yields:
$\frac{d\psi}{\psi}=-\frac{a_{0}}{\Gamma(T)}T^{\mu_{\psi}-1}(T_{c}-T)\,dT,$ (56)
which will be used to extract the scaling behavior of the order parameter near the critical temperature.
To obtain the desired critical scaling behavior, we now propose that the kinetic coefficient $\Gamma(T)$ vanishes near $T_{c}$, reflecting critical slowing down. Specifically, we take:
$\Gamma(T)=\frac{1}{\gamma}(T_{c}-T)^{2},$ (57)
where $\mu_{\psi}\in(0,1]$ is a conformable deformation parameter, and $\gamma$ is a dimensional constant.
Substituting back, we obtain the phenomenological differential equation:
$\frac{d\psi}{\psi}=-\gamma T^{\mu_{\psi}-1}(T_{c}-T)^{-1}\,dT$ (58)
This equation directly yields the critical behavior:
$\int\frac{d\psi}{\psi}$ $=-\gamma T^{\mu_{\psi}-1}\int\frac{dT}{(T_{c}-T)}.$ (59)
Integrating, we obtain
$\ln\psi(T)=\gamma T_{c}^{\mu-1}\ln|T-T_{c}|+\text{const.}$. (60)
So we can conclude that
$\psi(T)\sim(T_{c}-T)^{\beta(\mu)},\quad\text{with}\quad\beta(\mu)=\gamma T_{c}^{\mu_{\psi}-1},$ (61)
where we approximated $T^{\mu_{\psi}-1}\approx T_{c}^{\mu_{\psi}-1}$ near $T=T_{c}$. Thus, the critical exponent $\beta$ emerges naturally from a GL-type relaxation model with thermally modulated kinetics.
[…]
The critical exponent $\beta(\mu)$ can be viewed as a function encoding microscopic memory or fractal behavior in temperature evolution.
Deviations from the mean-field value occur when:
  • Thermal fluctuations near $T_{c}$ become significant (especially in $d<4$), requiring renormalization group treatment,
  • Nonlocal effects or memory are present, as modeled here via conformable derivatives,
  • The system is driven out of equilibrium, or experiences dissipation or long-time correlations.
These effects modify the scaling balance in the GL equation and yield $\beta>\tfrac{1}{2}$, consistent with experimental observations in real superconductors, which often exhibit critical behavior closer to the 3D XY universality class ($\beta\approx 0.35$).
VII.3 Emergence of Penetration Depth Scaling Exponent from Conformable Dynamics
The London penetration depth $\lambda_{L}(T)$ characterizes the distance over which magnetic fields decay inside a superconductor. Near the critical temperature $T_{c}$, it diverges as superconductivity breaks down. In analogy with critical phenomena, this divergence can be modeled by a deformed dynamical relation using conformable derivatives.
We begin with the assumption that $\lambda_{L}(T)$ obeys a conformable differential equation of the form:
$D_{T}^{(\mu)}\lambda_{L}(T)=-\gamma\,(T_{c}-T)^{-1}\lambda_{L}(T),$ (62)
where:
  • $D_{T}^{(\mu)}\lambda_{L}:=T^{1-\mu}\frac{d\lambda_{L}}{dT}$ is the conformable derivative,
  • $\mu\in(0,1]$ is the order of the deformation,
  • $\gamma$ is a positive scaling constant.
This equation states that the relative temperature derivative of $\lambda_{L}$, modulated by $T^{1-\mu}$, is proportional to the inverse distance from the critical point $T_{c}$. It encapsulates both the scaling nature of the divergence and possible memory/fractal effects through $\mu$.
We first substitute the definition of the conformable derivative:
$T^{1-\mu}\frac{d\lambda_{L}}{dT}=-\gamma(T_{c}-T)^{-1}\lambda_{L}(T)\quad\Rightarrow\quad\frac{d\lambda_{L}}{\lambda_{L}}=-\gamma\,T^{\mu-1}(T_{c}-T)^{-1}dT.$ (63)
Near the critical point $T\approx T_{c}$, we approximate $T^{\mu-1}\approx T_{c}^{\mu-1}$, yielding:
$\frac{d\lambda_{L}}{\lambda_{L}}\approx-\gamma\,T_{c}^{\mu-1}(T_{c}-T)^{-1}dT$ (64)
Integrate both sides of the equation above,
$\int\frac{d\lambda_{L}}{\lambda_{L}}=-\gamma T_{c}^{\mu-1}\int\frac{dT}{T-T_{c}}\quad\Rightarrow\quad\ln\lambda_{L}(T)=-\gamma T_{c}^{\mu-1}\ln|T-T_{c}|+\text{const.}$. (65)
And we obtain the scaling law given by
$\lambda_{L}(T)=B\,|T-T_{c}|^{-\alpha},\quad\text{with}\quad\alpha=\gamma T_{c}^{\mu-1}.$ (66)
This result shows that the penetration depth diverges as a power law near $T_{c}$, with an exponent that is not arbitrary but depends on:
the conformable order $\mu$, the critical temperature $T_{c}$ and the coupling constant $\gamma$.
Again, to model possible asymmetry in behavior above and below $T_{c}$, we generalize this to a piecewise form with different exponents, in a analogou way we do with $C_{V}$:
$\lambda_{L}(T)=\begin{cases}B_{1}\,|T-T_{c}|^{-\alpha_{1}},&T<T_{c},\\[4.0pt] B_{2}\,|T-T_{c}|^{-\alpha_{2}},&T\geq T_{c},\end{cases}\quad\text{where }\alpha_{i}=\gamma_{i}T_{c}^{\mu_{i}-1}$ (67)
Finally, to prevent divergence and ensure continuity near $T=T_{c}$, we introduce a regularization parameter $\epsilon>0$, resulting in the practical model used in the fit:
$\lambda_{L}(T)=B_{i}\left(|T-T_{c}|+\epsilon\right)^{-\alpha_{i}},\quad\text{for }i=\begin{cases}1,&T<T_{c}\\ 2,&T\geq T_{c}\end{cases}$
This regularized conformable model reproduces the observed smooth divergence and captures the asymmetric critical behavior seen in superconducting materials like niobium.
VII.3.1 Connection Between Penetration Depth and Order Parameter
The London penetration depth $\lambda_{L}(T)$ characterizes the distance over which an external magnetic field decays exponentially inside a superconductor. It provides a direct measure of the material’s ability to expel magnetic flux via the Meissner effect.
In Ginzburg–Landau theory, the superconducting state is described by a complex order parameter $\psi(T)$, where the squared magnitude $|\psi(T)|^{2}$ corresponds to the density of superconducting carriers. This connection leads to a fundamental relation between the order parameter and the electromagnetic response of the superconductor.
From the London equation, which relates the supercurrent $\mathbf{J}_{s}$ to the vector potential $\mathbf{A}$, we have:
×$\nabla\times\mathbf{J}_{s}=-\frac{n_{s}e^{2}}{m}\mathbf{B},$ (68)
where $n_{s}$ is the superconducting carrier density, $e$ is the elementary charge, $m$ is the effective mass, and $\mathbf{B}$ is the magnetic field. The London penetration depth is then given by:
$\lambda_{L}^{-2}(T)=\frac{\mu_{0}n_{s}(T)e^{2}}{m},$ (69)
where $\mu_{0}$ is the vacuum permeability.
Under the Ginzburg–Landau framework, the carrier density is proportional to the square of the order parameter: $n_{s}(T)\sim|\psi(T)|^{2}$. Substituting into the expression for $\lambda_{L}^{-2}(T)$ yields the key relation:
$\lambda_{L}^{-2}(T)\propto|\psi(T)|^{2}.$ (70)
This equation shows that as the temperature approaches the critical temperature $T_{c}$, the order parameter $\psi(T)\to 0$, and consequently $\lambda_{L}(T)\to\infty$. Physically, this reflects the loss of superconductivity and the full penetration of magnetic fields into the material.
Hence, the temperature dependence of $\lambda_{L}(T)$ provides a direct experimental probe of the order parameter’s behavior and is commonly used to extract critical exponents and validate theoretical models.
Near the critical temperature $T_{c}$, we have shown from the conformable Ginzburg-Landau equation that the superconducting order parameter behaves as:
$|\psi(T)|\sim\left(1-\frac{T}{T_{c}}\right)^{\beta},\qquad\text{for }T<T_{c},$ (71)
where the critical exponent $\beta$ depends on the conformable deformation parameter $\alpha$. This reduces to the standard mean-field behavior $\beta=\tfrac{1}{2}.$ When $\alpha=1$, recovering the classical Ginzburg-Landau result Tinkham (1996). By this way, near the critical temperature $T_{c}$, the superconducting order parameter vanishes as: where $\beta$ is the critical exponent for the order parameter in the Landau-Ginzburg theory or more general scaling theory.
Substituting Eq. (73) into Eq. (70) gives: In Ginzburg–Landau theory, the London penetration depth $\lambda_{L}(T)$ is directly related to the density of superconducting carriers. Specifically, it satisfies the relation:
$\lambda_{L}^{-2}(T)\propto|\psi(T)|^{2},$ (72)
where $\psi(T)$ is the order parameter describing the superconducting phase.
Near the critical temperature $T_{c}$, the order parameter vanishes as:
$|\psi(T)|\sim(T_{c}-T)^{\beta},\qquad\text{for }T<T_{c},$ (73)
which implies that:
$|\psi(T)|^{2}\sim(T_{c}-T)^{2\beta}.$ (74)
Substituting this into the expression for the penetration depth yields:
$\lambda_{L}^{-2}(T)\sim(T_{c}-T)^{2\beta}.$ (75)
Taking the inverse and square root, we obtain the leading behavior of $\lambda_{L}(T)$ as:
$\lambda_{L}(T)\sim(T_{c}-T)^{-\beta}.$ (76)
This result shows that the London penetration depth diverges as the temperature approaches $T_{c}$ from below. Therefore, the critical exponent governing its divergence is $\beta.$
This scaling relation is a direct consequence of the Ginzburg–Landau formalism and can be used to extract the critical exponent $\beta$ from experimental measurements of $\lambda_{L}(T)$ near the superconducting transition.
VII.4 Conformable Scaling of Specific Heat Near $T_{c}$: Heat Capacity Jump
In classical thermodynamics of phase transitions, the specific heat near the critical temperature $T_{c}$ typically scales as:
$C(T)\sim|T-T_{c}|^{-\alpha},$
where $\alpha$ is the critical exponent for the specific heat. In mean-field theory, $\alpha=0$, corresponding to a finite jump at $T_{c}$, but in more generalized or fluctuation-driven models, $\alpha$ can take nonzero values and describe a divergence.
To generalize this behavior within the conformable derivative framework, we have shown that the specific heat satisfies a deformed differential equation of the form of Eq. (9).
$C(T)\sim|T-T_{c}|^{-\alpha},\qquad\text{with}\qquad\alpha=\kappa T_{c}^{\mu_{C}-1}.$ (77)
This result shows that the critical exponent $\alpha$ is no longer a universal constant but depends on the deformation parameter $\mu_{C}$ and the critical temperature $T_{c}$. The conformable framework thus naturally introduces a generalized scaling law for the specific heat, while remaining analytically tractable and dimensionally consistent.
The generalized conformable framework enables direct fitting of experimental data from superconducting systems using critical scaling laws derived from deformed thermodynamic equations.
Specifically, the temperature dependence of key observables can be analyzed as follows: the order parameter $\psi(T)$ may be extracted from angle-resolved photoemission spectroscopy (ARPES), tunneling spectroscopy, or other coherence-sensitive measurements; the London penetration depth $\lambda_{L}(T)$ is accessible through microwave cavity perturbation, magnetic susceptibility, or muon spin rotation techniques; and the specific heat $C(T)$ is measured via calorimetry near the superconducting transition.
Each observable exhibits power-law scaling near the critical temperature $T_{c}$, governed by exponents that emerge naturally from the conformable derivative formulation.
This framework brings together different experimental observables under a unified algebraic structure, offering both analytical tractability and thermodynamic coherence. It serves as a phenomenological alternative to renormalization-group methods by capturing critical behavior through modified kinetic laws governed by the conformable derivative.
VII.5 Reinforcing justification of non-equilibrium deformed model
If solving the time-dependent or inhomogeneous problem, derivative terms must be kept and may dominate if $\psi(T)$ varies sharply.
But for static, near-equilibrium (mean-field) behavior, they contribute only to subleading corrections.
The Landau-based equation is an equilibrium condition: it comes from minimizing a free energy, assuming time-independence, homogeneity, and no fluctuations.
The conformable dynamic equation instead describes a non-equilibrium evolution, where the system may be relaxing toward equilibrium with scale-sensitive kinetics due to $\mu_{\psi}\neq 1.$
Therefore, the deformed dynamic equation gives a modified critical exponent, which reduces to $1/2$ only when $\mu_{\psi}=1$ and , i.e., when the dynamics are classical.
The exponent $\beta=1/2$ emerges from equilibrium minimization of the GL free energy, while $\beta=\gamma T_{C}^{\mu_{\psi}-1}$results from integrating the conformable dynamic evolution equation, which accounts for non-equilibrium scaling governed by the conformable derivative.
They agree only when the deformation vanishes.
VIII Numerical Fitting and Data Analysis
To test the predictive power of the conformable formalism, we apply it to experimental data for niobium’s superconducting transition.
VIII.1 Heat capacity of Niobium superconductor
To validate the conformable derivative model applied to superconducting transitions, we performed numerical curve fitting using experimental data.
Figure 1 shows the experimental heat capacity data for niobium near its superconducting transition, along with a conformable piecewise fit that captures the critical behavior on both sides of the transition temperature $T_{c}=8.700$ K. The fitted parameters are $B_{1}=0.411$, $\alpha_{1}=0.644$ for $T<T_{c}$, and $B_{2}=0.808$, $\alpha_{2}=0.567$ for $T>T_{c}$, with a smoothing factor $\epsilon=0.2928$ to ensure regularization near the divergence. As already commented, the model was regularized as
$C_{V}(T)=\begin{cases}B_{1}\left(|T-T_{c}|+\epsilon\right)^{-\alpha_{1}},&\text{for }T<T_{c}\\[6.0pt] B_{2}\left(|T-T_{c}|+\epsilon\right)^{-\alpha_{2}},&\text{for }T>T_{c}\end{cases}$ (78)
Figure 1: Heat capacity of niobium as a function of temperature near the superconducting transition, fitted using a piecewise conformable model. The fitted function follows the form $C_{V}(T)=B_{i}(|T-T_{c}|+\epsilon)^{-\alpha_{i}}$, with parameters: $B_{1}=0.411$, $\alpha_{1}=0.644$ (for $T<T_{c}$), and $B_{2}=0.808$, $\alpha_{2}=0.567$ (for $T>T_{c}$), where $\epsilon=0.2928$ and $T_{c}=8.700$ K. The black circles represent the digitized experimental data from Brown et al. (1953) Brown et al. (1953) .
VIII.2 Penetration Depth and Conformable Scaling
The temperature dependence of the London penetration depth $\lambda_{L}(T)$ provides key information about superconducting coherence and fluctuation behavior near the critical temperature $T_{c}$. Analogous to the divergence observed in specific heat, the penetration depth exhibits critical-like behavior and can be modeled using a conformable scaling framework.
We adopt the following conformable-inspired piecewise model:
$\lambda_{L}(T)=\begin{cases}B_{1}\left(|T-T_{c}|+\epsilon\right)^{-\alpha_{1}},&T<T_{c},\\[4.0pt] B_{2}\left(|T-T_{c}|+\epsilon\right)^{-\alpha_{2}},&T\geq T_{c},\end{cases}$ (79)
where $\epsilon>0$ is a regularization parameter that ensures finite behavior at the transition point.
This expression mirrors the scaling law:
$\lambda_{L}(T)\sim|T-T_{c}|^{-\alpha},$ (80)
with different exponents below and above the transition to capture the asymmetric behavior seen in experimental data. The conformable derivative framework again justifies this scaling law via the modified differential relation:
$D_{T}^{(\mu)}\lambda_{L}(T)=-\gamma(T-T_{c})^{-1}\lambda_{L}(T),\quad\Rightarrow\quad\lambda_{L}(T)\sim|T-T_{c}|^{-\gamma T_{c}^{\mu-1}}.$
[…]
The fitted parameters extracted from the data in Fig. 2 are:
$B_{1}=1.57,\quad\alpha_{1}=0.68,$
$B_{2}=5.01,\quad\alpha_{2}=0.98,$
$\epsilon=0.055,\quad T_{c}=9.78\leavevmode\nobreak\ \text{K}.$
Figure 2: Fit of the normalized London penetration depth $\lambda_{L}(T)/\lambda_{0}$ for niobium using a conformable piecewise model. The function $\lambda_{L}(T)=B_{i}(|T-T_{c}|+\epsilon)^{-\alpha_{i}}$, with $i=1$ for $T<T_{c}$ and $i=2$ for $T\geq T_{c}$, accounts for the asymmetry in critical behavior. The fitted parameters are: $B_{1}=1.57$, $\alpha_{1}=0.68$; $B_{2}=5.01$, $\alpha_{2}=0.98$; $\epsilon=0.055$, and $T_{c}=9.78\leavevmode\nobreak\ \mathrm{K}$. Data derived from: B. W. Maxfield and W. L. McLean, "Superconducting Penetration Depth of Niobium," Phys. Rev. 139, A1515–A1522 (1965), Fig.7. Digitized versus transformed variable and replotted as $\lambda(T)/\lambda_{0}$ vs $T$ Maxfield and McLean (1965).
VIII.3 Coherence Length from $H_{c2}(T)$ in Niobium
To validate the theoretical predictions, we turn to experimental data on niobium and extract the temperature-dependent coherence length using the upper critical field $H_{c2}(T)$.
VIII.3.1 Theoretical Background
In the Ginzburg–Landau (GL) framework, the upper critical magnetic field $H_{c2}(T)$ is related to the superconducting coherence length $\xi(T)$ by the expression:
$H_{c2}(T)=\frac{\Phi_{0}}{2\pi\xi^{2}(T)},$ (81)
where $\Phi_{0}=h/2e\approx 2.07\times 10^{-15}\,\text{Wb}$ is the magnetic flux quantum. This inverse-square relation permits the extraction of $\xi(T)$ from experimental measurements of $H_{c2}(T)$.
Near the critical temperature $T_{c}$, the coherence length diverges according to a critical power law:
$\xi(T)\sim(T_{c}-T)^{-\nu},$ (82)
This power-law behavior emerges naturally from the conformable Ginzburg-Landau framework when fractional spatial derivatives are incorporated. Similar approaches using fractional or stochastic generalizations of the GL equation have also been explored in dynamical settings Zhang et al. (2022). To see this, consider the modified GL free energy with a conformable derivative:
$|D_{x}^{(\mu)}\psi(x)|^{2}=x^{2(1-\mu)}\left(\frac{d\psi}{dx}\right)^{2},$ (83)
which replaces the standard gradient term $|\nabla\psi|^{2}$. The corresponding GL equation in one dimension becomes:
$a_{0}(T-T_{c})\psi+b|\psi|^{2}\psi-\frac{1}{2m}\left[x^{2(1-\mu)}\frac{d^{2}\psi}{dx^{2}}+(1-\mu)x^{1-2\mu}\frac{d\psi}{dx}\right]=0.$ (84)
Assuming a linearized form near $T_{c}$ and a trial solution $\psi(x)=\psi_{0}e^{-x/\xi}$, we compute:
$\frac{d\psi}{dx}$ $=-\frac{\psi_{0}}{\xi}e^{-x/\xi},$ (85)
$\frac{d^{2}\psi}{dx^{2}}$ $=\frac{\psi_{0}}{\xi^{2}}e^{-x/\xi}.$ (86)
Substituting into the kinetic term gives:
Kinetic $\text{Kinetic Term}\approx\frac{1}{2m}\left[\frac{x^{2(1-\mu)}}{\xi^{2}}+\frac{(1-\mu)x^{1-2\mu}}{\xi}\right]\psi(x).$ (87)
Balancing with the linear term $a_{0}(T-T_{c})\psi$, we find:
$a_{0}(T-T_{c})\sim\frac{1}{2m}\left[\frac{x^{2(1-\mu)}}{\xi^{2}}+\frac{(1-\mu)x^{1-2\mu}}{\xi}\right].$ (88)
In the limit $T\to T_{c}$ and assuming large $\xi$, the second term dominates, leading to:
$\xi\sim(T_{c}-T)^{-1},$ (89)
while in the opposite regime, the first term dominates and we recover:
$\xi^{2}\sim(T_{c}-T)^{-1}\Rightarrow\xi(T)\sim(T_{c}-T)^{-1/2}.$ (90)
Thus, the power-law divergence (82) is recovered with an effective exponent $\nu$ that may vary depending on which term dominates and the value of the conformable parameter $\mu$.
In fractional generalizations of GL theory, one may therefore consider replacing the usual spatial gradient term $|\nabla\psi|^{2}$ by this deformed expression involving the conformable derivative, leading to a modified scaling law for $\xi(T)$.
The experimental data for $H_{c2}(T)$ were extracted from Figure 4 of the classical work by Williamson Williamson (1970), which presents the behavior of the upper critical field for niobium. The experimental points (open circles) were digitized using the WebPlotDigitizer tool.
The original figure displays the data in a dimensionless form, with the vertical axis representing the ratio $-H_{c2}^{2}/\left(dH_{c2}/dt\right)_{t=1}$ as a function of the reduced temperature $t=T/T_{c}$. To reconstruct the actual values of $H_{c2}(T)$ in physical units (Tesla), we assumed a functional dependence near the critical transition given by:
$H_{c2}(T)=H_{c2}(0)\left(1-\frac{T}{T_{c}}\right)^{n},$ (91)
where $H_{c2}(0)$ denotes the extrapolated upper critical field at zero temperature, and $n$ is a fitting exponent.
The critical temperature was fixed at $T_{c}=9.2\,\text{K}$, consistent with values for pure niobium and in line with the parameters adopted in the literature Tinkham (1996).
After conversion, the data were used to fit the empirical model of Eq. (91), enabling the subsequent extraction of the temperature dependence of the coherence length $\xi(T)$ through the relation:
$\xi(T)=\sqrt{\frac{\Phi_{0}}{2\pi H_{c2}(T)}},$ (92)
where $\Phi_{0}=h/2e\approx 2.07\times 10^{-15}\,\text{Wb}$ is the magnetic flux quantum.
This procedure provides an experimental foundation to assess the validity of theoretical models for the temperature dependence of the coherence length.
Here, we claim that the use of a conformable derivative in the Ginzburg-Landau framework is motivated by the need to capture mesoscopic and spatially inhomogeneous effects that are not adequately described by the standard theory. In particular, real superconducting samples often exhibit intrinsic disorder, nonuniform pinning landscapes, or microstructural granularity, leading to effective nonlocal interactions and spatially varying coherence. The conformable derivative introduces a scale-dependent deformation of the spatial gradient, which can mimic the impact of fractal-like geometries or finite-size scaling in constrained domains. This modification allows for a more flexible description of the kinetic term and leads to a generalized coherence length scaling, making the theory more compatible with experimental deviations from classical mean-field behavior.
VIII.3.2 Results and Analysis
Figure 3 displays the digitized data and the fitted model. The excellent agreement indicates the suitability of the power-law model in this temperature range.
Figure 3: Extracted coherence length $\xi(T)$ (blue dots) and fitted model (red line) using $\xi(T)=A(T_{c}-T)^{-\nu}$ with $\nu=0.329;$$A=40.76$. Experimental data adapted from Williamson et al. Williamson (1970).
The coherence length $\xi(T)$ describes the spatial extent over which the superconducting order parameter $\psi$ varies appreciably. As $T\to T_{c}$, $\xi(T)\to\infty$, and $A$ sets the overall magnitude of $\xi$ for temperatures below $T_{c}$.
The coefficient $A$ encodes material-specific properties, including microscopic interactions, effective mass of Cooper pairs, and strength of the pairing potential.
The extracted critical exponent $\nu\approx 0.329$ is notably smaller than the classical Ginzburg–Landau (GL) mean-field prediction of $\nu=1/2$. Several factors may contribute to this deviation:
  • Limited resolution or experimental uncertainty in the high-temperature region ($T\to T_{c}$);
  • Residual disorder or impurity-induced scattering effects not incorporated into the ideal GL framework;
  • Memory effects—either thermal or spatial—captured through conformable (fractional-order) derivatives;
  • Effective averaging over anisotropic coherence lengths in polycrystalline or textured samples.
From the expression:
$\xi(T)=A(T_{c}-T)^{-\nu},$ (93)
we deduce the units of $A$ by dimensional analysis. Since $\xi(T)$ is in nanometers (nm), $(T_{c}-T)$ is in Kelvin (K), and $\nu$ is dimensionless, then $A$ must have units of nm·Kν.
Given the fitted exponent $\nu\approx 0.329$ and $A=40.76$, this means:
[…]
To express $A$ in SI units (meters), we convert:
$A_{\text{SI}}=40.76\times 10^{-9}\ \text{m}\cdot\text{K}^{0.329}=4.076\times 10^{-8}\ \text{m}\cdot\text{K}^{0.329}.$ (95)
Such a reduced exponent is qualitatively compatible with generalized GL models that incorporate non-integer-order differential structures.
For clean elemental superconductors like niobium, coherence lengths on the order of 30–40 nm are typical at intermediate temperatures. The value $A=40.76nm\cdot K^{\nu}$ is thus physically reasonable and consistent with known properties of niobium. It represents the critical amplitude of the coherence length near the critical temperature $T_{c}$. Its physical meaning and units depend on how the data was scaled in the fitting process.
The value of $A$ determines how quickly the coherence length diverges as temperature approaches $T_{c}$.
A larger $A$ implies that $\xi(T)$ becomes large faster as $T\to T_{c}$, for a fixed $\nu$. This can be influenced by material purity, anisotropy or also electron-phonon coupling strength.
VIII.3.3 Comparison to Literature
For clean elemental superconductors like niobium, coherence lengths on the order of 30–40 nm are typical at intermediate temperatures. The value $A=40.76$ nm.Kν is thus physically reasonable and consistent with known properties of niobium.
We conclude by summarizing the key results and highlighting how the conformable framework opens new directions in the study of critical dynamics, both in theory and experiment.
IX Conclusions and Outlook
[…]
In particular, the observed deviations from mean-field exponents are explained as consequences of conformable deformation, capturing finite-size effects, spatial granularity, and critical slowing down. The flexibility of the deformation parameter enables modeling of crossover behaviors between distinct universality classes, potentially capturing nontrivial fixed-point trajectories. The conformable approach thus provides a meaningful interpolation between classical mean-field theory and more general statistical frameworks, such as nonextensive thermodynamics.
A significant advantage of the conformable framework lies in its analytical tractability. Unlike renormalization group methods, which often rely on asymptotic expansions and perturbative schemes, the conformable model yields closed-form expressions that can be directly fitted to experimental data. This was demonstrated through application to superconducting phase transitions in niobium, where the model successfully captured the asymmetric scaling of specific heat and London penetration depth near the critical temperature. The extracted exponents are consistent with deviations from mean-field theory and suggest the presence of mesoscopic effects, spatial inhomogeneities, and memory-driven dynamics.
Its geometric interpretation, linked to deformations in the thermal or spatial metric Weberszpil and Helayël-Neto (2015); Rosa and Weberszpil (2018); Has et al. (2024), offers insight into how fractal geometry and nonlocal responses might influence critical behavior. In parallel, conformable symmetry approaches have also found applications in nuclear structure and critical point symmetriesRaissi-Rad et al. (2021), illustrating the broader relevance of such deformations across diverse physical systems
The framework is general and extensible. It can be adapted to a broad range of critical systems, including percolation thresholds, spin glasses, disordered media, and quantum phase transitions. Its generality makes it a valuable tool for exploring universality in complex systems, particularly in non-equilibrium or nonextensive regimes.
Looking ahead, the conformable formalism holds promise for broader applications in systems where classical scaling laws break down due to disorder, finite-size constraints, or complex topology or entropy-driven irreversibility Sotolongo-Costa and Weberszpil (2021). Similar approaches using structural or deformed derivatives have been proposed to model ultraslow diffusion in complex systems Xu et al. (2017). It may also serve as a foundation for modeling anomalous transport, quantum phase transitions, or nonequilibrium critical dynamics. Further investigation into its connection with renormalization group flows, scaling functions, and anomalous dimensions could deepen our understanding of universality in deformed or emergent settings.
Beyond phenomenological applications, several open questions remain in the rigorous mathematical formulation of the conformable framework. These include the development of a well-defined operator semigroup structure, the spectral theory of deformed differential operators, and the formulation of variational principles compatible with conformable dynamics Godinho et al. (2020). Addressing these foundational issues may reveal deeper connections between the conformable approach and the broader structure of mathematical physics.
Acknowledgments:
One of the authors, José Weberszpil, wishes to express their gratitude to FAPERJ, APQ1, for the partial financial support.
Declarations
Funding: The author received partial financial support from FAPERJ/APQ1.
Conflicts of interest/Competing interests: The author declares no conflicts of interest or competing interests related to this work.
Appendix A Dimensional Consistency Analysis
The mathematical structure of the conformable derivative suggests deeper geometric meaning. In this section, we explore interpretations based on fractal metrics and spatial deformation.
In this section, we analyze the dimensional consistency of the primary equations and critical exponents derived using the conformable derivative framework. This step is essential to ensure the physical viability of the proposed model.
A.1 Conformable Derivative Units
The conformable derivative with respect to temperature is defined as:
$D_{T}^{(\mu)}f(T)=T^{1-\mu}\frac{df}{dT}$ (96)
Letting $[T]=\text{K}$ (Kelvin), we note that $\frac{df}{dT}$ has units of $[f]/\text{K}$. Therefore:
$[D_{T}^{(\mu)}f]=[f]\cdot\text{K}^{-\mu}$ (97)
This confirms that the conformable derivative scales as expected, preserving consistent units when applied to temperature-dependent quantities.
A.2 Critical Exponent Dimensionality
The critical exponents $\alpha,\beta,\gamma,\nu$ are defined through relations such as:
$\alpha=\kappa T_{c}^{\mu_{C}-1},\quad\beta=\gamma T_{c}^{\mu_{M}-1},\quad\gamma=\lambda T_{c}^{\mu_{\chi}-1},\quad\nu=\rho T_{c}^{\mu_{\xi}-1}$ (98)
Since exponents like $\alpha$ appear in power-law expressions of the form $C(T)\sim|T-T_{c}|^{-\alpha}$, they must be dimensionless. For this to hold, $\kappa$ must have units of $\text{K}^{1-\mu_{C}}$, and similarly for the other coefficients. Hence, all critical exponents are indeed dimensionless, satisfying the necessary conditions for physical consistency.
A.3 Modified Ginzburg–Landau Equation
The modified Ginzburg–Landau equation derived using the conformable kinetic term is:
$a(T)\psi+b|\psi|^{2}\psi-\frac{1}{2m}T^{2(1-\alpha)}\frac{d^{2}\psi}{dT^{2}}-\frac{1-\alpha}{m}T^{1-2\alpha}\frac{d\psi}{dT}=0$ (99)
Let $\psi$ have units $[\psi]=\text{K}^{-1/2}$. Then:
  • $a(T)=a_{0}(T-T_{c})$, with $[a_{0}]=\text{K}^{-2}$
  • $b$ must have units to make $b|\psi|^{2}\psi$ have the same units as $a(T)\psi$: $\text{K}^{-5/2}$
  • Kinetic terms involve second and first derivatives, and with the conformable powers of $T$, they retain the same dimension as $\psi/\text{K}$, ensuring consistency.
Thus, all terms in the equation have matching units.
A.4 Heat Capacity Scaling
Heat capacity scales as:
$C(T)\sim|T-T_{c}|^{-\alpha},\quad[C]=\frac{E}{T}$ (100)
For dimensional consistency, $\alpha$ must be dimensionless, which is satisfied given that $\alpha=\kappa T_{c}^{\mu_{C}-1}$.
A.5 Penetration Depth and Order Parameter
The London penetration depth scales as:
$\lambda_{L}(T)\sim(T_{c}-T)^{-\zeta},\quad\zeta=\beta/2$ (101)
Assuming $[\lambda_{L}]=\text{length}$, and that $\zeta$ is dimensionless, the units are correctly preserved.
In summary, all core expressions and critical exponent definitions derived from the conformable derivative framework are dimensionally consistent. The use of temperature scaling through powers of $T$ in the deformed derivatives introduces no contradictions and supports the model’s physical credibility.
This ensures the robustness of the theoretical framework and its applicability to thermodynamical systems such as superconducting transitions.
Appendix B Variational Derivation of the Modified GL Equation
We start from the conformable free energy functional:
[…]
The conformable derivative with respect to temperature is:
$T_{\alpha}\psi(T)=T^{1-\alpha}\frac{d\psi}{dT}$ (103)
Then, the kinetic term becomes:
$\left|T_{\alpha}\psi(T)\right|^{2}=T^{2(1-\alpha)}\left|\frac{d\psi}{dT}\right|^{2}$ (104)
The full functional reads:
[…]
To minimize the free energy, we compute the variation with respect to $\psi^{*}(T)$:
[…]
Using integration by parts and assuming vanishing boundary terms:
[…]
Therefore, setting $\delta\mathcal{F}_{\alpha}=0$, the Euler–Lagrange equation becomes:
$a(T)\psi+b|\psi|^{2}\psi-\frac{1}{2m}\frac{d}{dT}\left(T^{2(1-\alpha)}\frac{d\psi}{dT}\right)=0$ (108)
Expanding the total derivative, we obtain:
$a(T)\psi+b|\psi|^{2}\psi-\frac{1}{2m}T^{2(1-\alpha)}\frac{d^{2}\psi}{dT^{2}}-\frac{(1-\alpha)}{m}T^{1-2\alpha}\frac{d\psi}{dT}=0$ (109)
B.1 Final Form of the Equation in Powers of $(1-\frac{T}{T_{c}})$
Substituting the ansatz
$\psi(T)=\psi_{0}\left(1-\frac{T}{T_{c}}\right)^{\beta}$ (110)
into the full equation:
$a_{0}(T-T_{c})\psi_{0}\left(1-\frac{T}{T_{c}}\right)^{\beta}+b\psi_{0}^{3}\left(1-\frac{T}{T_{c}}\right)^{3\beta}$
$\quad-\frac{1}{2m}T^{2(1-\alpha)}\frac{\beta(\beta-1)\psi_{0}}{T_{c}^{2}}\left(1-\frac{T}{T_{c}}\right)^{\beta-2}$
$\quad+\frac{(1-\alpha)}{m}T^{1-2\alpha}\frac{\beta\psi_{0}}{T_{c}}\left(1-\frac{T}{T_{c}}\right)^{\beta-1}=0$ (111)
Matching powers of $\left(1-\frac{T}{T_{c}}\right)$, we recover $\beta=\frac{1}{2}$ as the critical exponent.
Now, we will analyze the behavior of $\psi(T)$ near the critical temperature $T_{c}$. Consider again the nonlinear temperature-deformed Ginzburg–Landau-type equation:
$a(T)\psi+b|\psi|^{2}\psi-\frac{1}{2m}T^{2(1-\alpha)}\frac{d^{2}\psi}{dT^{2}}-\frac{1-\alpha}{m}T^{1-2\alpha}\frac{d\psi}{dT}=0$ (112)
with $a(T)=a_{0}(T-T_{c})$.
Assume again the ansatz:
$\psi(T)=\psi_{0}\left(1-\frac{T}{T_{c}}\right)^{\beta}=\psi_{0}\varepsilon^{\beta},\quad\text{where }\varepsilon=1-\frac{T}{T_{c}}$ (113)
Then, deriving in relation to T, gives,
$\frac{d\varepsilon}{dT}=-\frac{1}{T_{c}}$ (114)
$\frac{d\psi}{dT}$ $=-\frac{\beta\psi_{0}}{T_{c}}\varepsilon^{\beta-1}$ (115)
$\frac{d^{2}\psi}{dT^{2}}$ $=\frac{\beta(\beta-1)\psi_{0}}{T_{c}^{2}}\varepsilon^{\beta-2}$ (116)
In the sequence, substituting the derivatives into the Equation, each term becomes:
  • Linear term:
    $a_{0}(T-T_{c})\psi=-a_{0}T_{c}\psi_{0}\varepsilon^{\beta+1}$ (117)
  • Nonlinear term:
    $b|\psi|^{2}\psi=b\psi_{0}^{3}\varepsilon^{3\beta}$ (118)
  • Second derivative term:
    $-\frac{1}{2m}T^{2(1-\alpha)}\cdot\frac{\beta(\beta-1)\psi_{0}}{T_{c}^{2}}\varepsilon^{\beta-2}$ (119)
  • First derivative term:
    $\frac{(1-\alpha)\beta\psi_{0}}{mT_{c}}T^{1-2\alpha}\varepsilon^{\beta-1}$ (120)
Let’s try to balance the derivative terms with the nonlinear term and compare $\varepsilon^{\beta-2}$, $\varepsilon^{\beta-1}$ and $\varepsilon^{3\beta}.$
We seek a consistent balance. Suppose that the nonlinear term $\varepsilon^{3\beta}$ and the linear term $\varepsilon^{\beta+1}$dominate. Then equating powers: $3\beta=\beta+1,$that implies in $\beta=1/2.$
This is consistent with Landau theory, and a typical critical exponent in second-order phase transitions.
Then:
$\varepsilon^{\beta-2}=\varepsilon^{-3/2}$ (121)
$\varepsilon^{\beta-1}=\varepsilon^{-1/2}$
$\varepsilon^{3\beta}=\varepsilon^{3/2}$
$\varepsilon^{\beta+1}=\varepsilon^{3/2}$
So derivative terms diverge as $\varepsilon\to 0$, but despite their divergence, derivative terms describe kinetic effects (gradients of $\psi$). Near $T=T_{c}$, the order parameter becomes small and smooth: $\psi\to 0$, and its variation is slow and finnaly the mean-field and equilibrium theories focus on minimizing free energy — gradient terms are subleading.
Therefore, we match:
$a(T)\psi+b\psi^{3}=0\Rightarrow\psi=\pm\sqrt{-\frac{a(T)}{b}}$ (122)
giving again:
$\psi(T)=\psi_{0}\left(1-\frac{T}{T_{c}}\right)^{1/2}$ (123)
The derivative terms contribute to corrections to scaling or dynamical behavior, but not to the leading-order critical exponent.
So, we have demonstrated that the ansatz, 110, satisfies the nonlinear equation near $T=T_{c}$, with leading order balance provided by the algebraic terms:
$a(T)\psi+b\psi^{3}=0\Rightarrow\boxed{\beta=\frac{1}{2}}$ (124)
This matches classical Landau theory and remains valid under conformable kinetic corrections. Also, this represents a static equilibrium solution where the derivatives are zero (i.e., long-time equilibrium). It is important to note that our model gives that, under equilibrium, homogeneous conditions, and negligible fluctuations, $\beta=\frac{1}{2}.$
In summary, we have consided the nonlinear differential equation:
$a_{0}(T-T_{c})\psi+b\psi^{3}-\frac{1}{2m}T^{2(1-\alpha)}\psi^{\prime\prime}-\frac{1-\alpha}{m}T^{1-2\alpha}\psi^{\prime}=0$ (125)
Using the ansatz:
$\psi(T)=\psi_{0}\varepsilon^{\beta},\quad\text{with }\varepsilon=1-\frac{T}{T_{c}}$ (126)
and expanding near $\varepsilon\to 0^{+}$, we obtain:
Term Expression Leading Power of $\varepsilon$
Linear term $a_{0}(T-T_{c})\psi=-a_{0}T_{c}\psi_{0}\varepsilon^{\beta+1}$ $\varepsilon^{\beta+1}$
Nonlinear term $b\psi^{3}=b\psi_{0}^{3}\varepsilon^{3\beta}$ $\varepsilon^{3\beta}$
First derivative term $\frac{(1-\alpha)\beta\psi_{0}}{mT_{c}}T^{1-2\alpha}\varepsilon^{\beta-1}$ $\varepsilon^{\beta-1}$
Second derivative term $-\frac{1}{2m}T^{2(1-\alpha)}\cdot\frac{\beta(\beta-1)\psi_{0}}{T_{c}^{2}}\varepsilon^{\beta-2}$ $\varepsilon^{\beta-2}$
Table 1: Detailed expressions and their leading powers of $\varepsilon$ for each term in the equation.
As $\varepsilon\to 0$, any power $\varepsilon^{p}$ with $p<0$ diverges and would seem dominant—unless its coefficient is zero or negligible.
Substitute $\beta=\frac{1}{2}$, then:
$\varepsilon^{\beta-2}=\varepsilon^{-3/2},\quad\varepsilon^{\beta-1}=\varepsilon^{-1/2},\quad\varepsilon^{3\beta}=\varepsilon^{3/2},\quad\varepsilon^{\beta+1}=\varepsilon^{3/2}$
Clearly, $\varepsilon^{-3/2}$ and $\varepsilon^{-1/2}$ diverge more than $\varepsilon^{3/2}$, so they may seem to dominate.
In perturbative and scaling analyses:
  • We seek a consistent leading-order balance, not a match of all terms.
  • We match the dominant algebraic (non-derivative) terms: the linear and nonlinear terms.
  • This approach is justified because near $T\to T_{c}$, $\psi\to 0$, and its gradients $\psi^{\prime},\psi^{\prime\prime}$ become negligible — especially in spatially or thermally uniform systems.
Thus:
$a(T)\psi+b\psi^{3}=0\quad\Rightarrow\quad\psi(T)=\psi_{0}\left(1-\frac{T}{T_{c}}\right)^{1/2}$ (127)
Important: Derivative terms describe response to inhomogeneities (e.g., $\nabla\psi$), not the bulk equilibrium.
Appendix C Recovery of from the Relaxation Equation in the Undeformed Limit
We now show that the classical critical exponent $\beta=\tfrac{1}{2}$ naturally emerges from the conformable relaxation equation in the undeformed equilibrium limit. Consider the generalized relaxation model:
$\Gamma(T)\,T^{1-\mu_{\psi}}\frac{d\psi}{dT}=-a_{0}(T_{c}-T)\psi-b|\psi|^{2}\psi,$ (128)
where $\Gamma(T)$ is a temperature-dependent kinetic coefficient, $\mu_{\psi}\in(0,1]$ is the conformable deformation parameter, and the right-hand side arises from the variational derivative of the standard Ginzburg–Landau free energy.
In the undeformed limit, we take:
$\mu_{\psi}=1,\quad\text{so that }T^{1-\mu_{\psi}}=1,$ (129)
and for simplicity assume that the kinetic coefficient is constant near $T_{c}$, i.e., $\Gamma(T)\approx\Gamma_{0}$. Equation (128) becomes:
$\Gamma_{0}\frac{d\psi}{dT}=-a_{0}(T_{c}-T)\psi-b|\psi|^{2}\psi.$ (130)
We now look for a solution near the critical point of the form:
$\psi(T)=\psi_{0}(T_{c}-T)^{\beta},$ (131)
valid for $T<T_{c}$, with $\psi_{0}$ a constant amplitude to be determined. Computing the derivative:
$\frac{d\psi}{dT}=-\beta\psi_{0}(T_{c}-T)^{\beta-1}.$ (132)
Substituting into Eq. (130), we obtain:
$-\Gamma_{0}\beta\psi_{0}(T_{c}-T)^{\beta-1}=-a_{0}(T_{c}-T)\psi_{0}(T_{c}-T)^{\beta}-b\psi_{0}^{3}(T_{c}-T)^{3\beta}.$ (133)
Divide both sides by $\psi_{0}(T_{c}-T)^{\beta-1}$ (assuming $\psi_{0}\neq 0$ and $T<T_{c}$):
$-\Gamma_{0}\beta=-a_{0}(T_{c}-T)^{2}-b\psi_{0}^{2}(T_{c}-T)^{2\beta+1}.$ (134)
Now examine the leading-order behavior as $T\to T_{c}^{-}$: - The term $(T_{c}-T)^{2}$ dominates, - The nonlinear term $\propto(T_{c}-T)^{2\beta+1}$ is subleading if $2\beta+1>2\Rightarrow\beta>\tfrac{1}{2}$, or comparable if $\beta=\tfrac{1}{2}$.
So, assuming the dominant balance occurs between:
$-\Gamma_{0}\beta\approx-a_{0}(T_{c}-T)^{2},$ (135)
we see that this leads to a contradiction: the left-hand side is constant, the right-hand side vanishes as $T\to T_{c}$. Therefore, this balance fails.
Instead, assume the dominant balance occurs between the first derivative terma nd the nonlinear cubic term:
$-\Gamma_{0}\beta\psi_{0}(T_{c}-T)^{\beta-1}\approx-b\psi_{0}^{3}(T_{c}-T)^{3\beta}.$ (136)
Cancel signs and factors of $\psi_{0}$:
$\Gamma_{0}\beta(T_{c}-T)^{\beta-1}=b\psi_{0}^{2}(T_{c}-T)^{3\beta}.$ (137)
Divide both sides by $(T_{c}-T)^{\beta-1}$:
$\Gamma_{0}\beta=b\psi_{0}^{2}(T_{c}-T)^{2\beta+1}.$ (138)
To keep both sides finite and nonzero as $T\to T_{c}$, we must have:
$2\beta+1=0\quad\Rightarrow\quad\beta=-\tfrac{1}{2},$ (139)
which is unphysical.
So we try instead to match the linear and nonlinear terms (equivalent to the adiabatic limit):
$a_{0}(T_{c}-T)\psi\sim b\psi^{3}\Rightarrow a_{0}(T_{c}-T)\sim b\psi^{2}.$ (140)
Solving for $\psi(T)$:
$\psi^{2}(T)\sim\frac{a_{0}}{b}(T_{c}-T),\quad\Rightarrow\quad\psi(T)\sim\sqrt{\frac{a_{0}}{b}}(T_{c}-T)^{1/2}.$
Thus, the leading-order balance occurs not with the derivative term, but with the free-energy terms, confirming that in the long-time (quasi-equilibrium) limit, the stationary solution satisfies:
$\beta=\tfrac{1}{2}.$ (141)
This demonstrates that even when starting from the generalized relaxation equation, the classical mean-field exponent $\beta=\tfrac{1}{2}$ is recovered in the undeformed limit $\mu_{\psi}=1$, under equilibrium assumptions. The derivative terms decay near $T_{c}$, and the critical scaling is governed by the balance of potential terms in the Ginzburg–Landau free energy.
C.1 Higher-Order Corrections
Derivative terms can be treated as perturbations (next-to-leading order). They modify the scaling only slightly and do not affect the critical exponent at leading order.
  • $\varepsilon^{\beta-1}$ and $\varepsilon^{\beta-2}$ diverge as $\varepsilon\to 0$, but are derivative-based.
  • Near $T_{c}$, $\psi$ is small and smooth, so derivatives are small.
  • Dominant balance is between $a(T)\psi$ and $b\psi^{3}$, yielding again $\beta=\frac{1}{2}.$
  • Derivative terms are corrections to this leading-order behavior.