### v2

# Conformable scaling and critical phenomena: A unified framework for phase transitions

Abstract
We investigate the application of conformable derivatives to model critical phenomena near continuous phase transition points. By incorporating a deformation parameter into the differential structure, we derive unified expressions for thermodynamic observables such as heat capacity, magnetization, susceptibility, and coherence length, each exhibiting a power-law behavior near the critical temperature. The conformable derivative framework naturally embeds scale invariance and critical slowing down into the dynamics without resorting to fully nonlocal fractional calculus. Modified Ginzburg–Landau equations are constructed to model superconducting transitions, leading to analytical expressions for the order parameter and the London penetration depth. Experimental data from niobium confirm the model’s applicability, showing excellent fits and capturing asymmetric scaling behavior around the critical point. This work offers a bridge between classical mean-field theory and generalized scaling frameworks, with implications for both theoretical modeling and experimental analysis.
Keywords: Conformable derivative; Modified Ginzburg-Landau equation;Critical phenomena; Critical scaling; Continuous phase transition; Thermodynamic consistency; Emergent dynamics; Power-law behavior.
I Introduction
Critical phenomena play a central role in the understanding of phase transitions, where physical observables such as the heat capacity, magnetization, and susceptibility exhibit a non-analytical behavior characterized by power-law divergences near a critical temperature $T_{c}$ Landau and Lifshitz (1978); Goldenfeld (1992); Stanley (1971). These behaviors are traditionally described through scaling hypotheses and renormalization group (RG) analysis, which explain the emergence of universality across diverse systems. These phenomena range from magnetic materials and superconductors to liquid-gas transitions and even biological systems Goldenfeld (1992); Stanley (1971). Near critical points, physical systems exhibit remarkable scaling behavior characterized by power-law divergences of thermodynamic quantities, with critical exponents that depend only on fundamental symmetries and spatial dimensionality rather than microscopic details Wilson (1971); Fisher and Barber (1972). This universality principle, established through decades of theoretical and experimental work, forms the cornerstone of the modern theory of critical phenomena.
The mathematical description of critical behavior has evolved significantly since the early phenomenological approaches of Landau and Ginzburg Landau (1965); Ginzburg and Landau (1965). The subsequent development of Ginzburg-Landau theory, particularly as systematized by de Gennes de Gennes (1966), provided a robust framework for understanding superconducting phase transitions and scaling behavior near critical points. The development of renormalization group theory by Wilson and others Wilson (1971); Wilson and Kogut (1974) provided deep insights into the origin of universality and scaling laws, while field-theoretic methods enabled precise calculations of critical exponents Zinn-Justin (2021). These advances established that critical phenomena emerge from the competition between thermal fluctuations and ordering tendencies, with correlation lengths diverging as $\xi\sim|T-T_{c}|^{-\nu}$ and other observables following characteristic power-laws.
Despite these theoretical successes, many experimental systems exhibit deviations from ideal critical behavior due to finite-size effects, quenched disorder, non-equilibrium conditions, and other realistic complications Privman (1990); Harris (1974); Hohenberg and Halperin (1977b). There remains a need for alternative frameworks that provide analytical tractability, flexible modeling, and connections to generalized thermodynamic formalisms. This is particularly relevant in systems where microscopic details are poorly known, long-range interactions are present, or nonlocal and memory effects become significant. Real materials often deviate from ideal critical behavior due to quenched disorder, finite-size effects, and non-equilibrium conditions. These deviations often manifest as modified scaling laws, anomalous relaxation dynamics, and effective critical exponents that differ from theoretical predictions. Understanding and modeling such complex critical behavior remains an active area of research with important implications for materials science, statistical physics, and complex systems.
In this work, we explore a deformation-based approach to critical phenomena using conformable derivatives Weberszpil and Helayël-Neto (2015, 2016); Weberszpil and Sotolongo-Costa (2017); Weberszpil and Chen (2017); Sotolongo-Costa and Weberszpil (2021); Rosa and Weberszpil (2018); Godinho et al. (2020); Khalil et al. (2014), a form of generalized calculus that introduces a deformation parameter $\mu$ into the derivative operator:
$D_{T}^{(\mu)}f(T):=T^{1-\mu}\frac{df}{dT}.$ (1)
This operator naturally yields temperature-dependent scaling behavior and introduces power-law features without requiring full fractional integration or nonlocal convolution terms. It thus offers an intermediate practical framework settled between classical and fractional derivatives; and retains analytical simplicity while capturing essential scaling features.
The conformable derivative framework addresses specific limitations in existing approaches to critical phenomena. Real materials exhibit: (1) Spatial inhomogeneities leading to effective fractal geometries with non-integer dimensions; (2) Temporal memory effects from quenched disorder or slow relaxation processes; and (3) Finite-size constraints that modify critical scaling. Each of these physical mechanisms naturally leads to temperature-dependent weighting factors of the form $T^{1-\mu}$, providing direct physical justification for definition (1). Unlike purely phenomenological approaches, the conformable-derivative framework connects the deformation parameter $\mu$ to measurable physical quantities: fractal dimensions in porous superconductors, anomalous diffusion exponents in disordered systems, and finite-size scaling exponents in thin films.
We note that while Eq. (1) might appear to be merely a change of variables $T\to T^{\prime}$ with a power-law transformation, the physical motivation and mathematical implementation differ fundamentally from simple rescaling. Namely, the conformable derivative (1) introduces a temperature-dependent weighting that reflects a physical process rather than mathematical convenience. These process aspects include: (i) anomalous diffusion and memory effects in disordered systems (Metzler and Klafter, 2000, 2004; Klafter and Sokolov, 2011; Sokolov, 2012; Metzler et al., 2014), (ii) finite-size scaling in confined geometries (Fisher and Barber, 1972; Binder, 1981; Privman, 1990), and (iii) non-equilibrium relaxation dynamics near criticality (Hohenberg and Halperin, 1977b; Bray, 1994; Godrèche and Luck, 2000). Unlike a simple coordinate transformation, the deformation parameter $\mu$ emerges from the underlying physics and connects to measurable quantities such as fractal dimensions and effective transport exponents. Furthermore, the critical exponents derived here are not invariant under the transformation because the physical interpretation changes. The parameter $\mu$ captures deviations from mean-field behavior that would otherwise require complex renormalization group calculations or phenomenological fitting.
Formally, $D_{T}^{(\mu)}f=T^{1-\mu}\,df/dT$ can be rewritten as a derivative with respect to $T^{\prime}=\tfrac{T^{\mu}}{\mu}$ only if all coefficients are constant or trivial functions of $T^{\prime}$. In our case this is not true for two reasons: (i) the kinetic coefficient $\Gamma(T)\sim|T-T_{c}|^{z\nu}$ retains its critical singularity under the transformation, preserving nontrivial scaling; (ii) in the conformable Ginzburg-Landau (GL) functional, the gradient term acquires a weight $T^{2(1-\alpha)}$, producing a non-uniform thermal metric, as we will see. After reparametrization, Jacobian factors survive and prevent a reduction to a standard GL form. Thus the conformable framework encodes a dynamical and variational structure that is not removed by a change of variables.
Here and throughout this work, the symbol $\sim$” denotes asymptotic proportionality: for two quantities $A$ and $B$, the relation $A\sim B$ means that their ratio tends to a finite, non-zero constant as the critical point is approached,
$\lim_{T\to T_{c}}\frac{A}{B}=C,\qquad 0<C<\infty.$
In other words, $A$ and $B$ share the same leading power-law behavior near the singularity, while multiplicative prefactors may differ.
We apply this formalism to describe thermodynamic observables near continuous phase transitions, including heat capacity, magnetization, susceptibility, and correlation length. In each case, we derive the critical exponents as functions of the deformation parameter and temperature, leading to expressions compatible with known mean-field results. Furthermore, we extend the method to superconducting phase transitions by constructing a modified Ginzburg–Landau (MGL) equation with conformable kinetic terms. This allows us to analyze the temperature dependence of the superconducting order parameter, London penetration depth, and heat capacity within the same mathematical framework.
We argue that with this conformable-derivative approach it is possible to obtain a unified formalism for deriving all major critical exponents with analytical expressions that match experimental data with high fidelity. This formalism thus represents a bridge between classical mean-field theory and generalized (e.g., Tsallis) statistical mechanics—and also a minimal extension of known differential equations without resorting to nonlocal fractional operators.
By maintaining dimensional consistency and compatibility with equilibrium thermodynamics, the conformable-derivative framework provides a useful and versatile tool for exploring critical dynamics in both classical and quantum systems. The results presented here demonstrate the capacity of this formalism to reproduce known scaling laws while offering new insights into generalized scaling behavior near criticality.
We derive unified expressions for critical exponents in terms of conformable parameters that emerge from these physical mechanisms, demonstrating that the approach extends rather than contradicts renormalization group theory. The formalism is applied to magnetic phase transitions, superconducting transitions, and other critical phenomena, with specific predictions for experimental verification.
Critical dynamics, relaxation forms, and ergodicity
Dynamic scaling and critical slowing down.
Near $T_{c}$ the characteristic relaxation time diverges as
$\tau\sim\xi^{\,z}\sim|T-T_{c}|^{-z\nu},$ (2)
with dynamic exponent $z$ and correlation-length exponent $\nu$. Time-dependent correlations obey the scaling form
$C_{A}(t;\,T)\equiv\langle A(t)A(0)\rangle_{c}\;=\;t^{-\,\lambda_{C}/z}\;\mathcal{F}\!\left(\frac{t}{\tau}\right),$ (3)
where $\lambda_{C}$ is a (model-dependent) autocorrelation exponent and $\mathcal{F}(x)$ is a universal scaling function. Far from criticality ($t\!\ll\!\tau$) one typically observes exponential decay; at $T\simeq T_{c}$ ($\tau\!\to\!\infty$) power-law tails dominate:
$C_{A}(t)\;\propto\;t^{-\,\lambda_{C}/z}\qquad(T=T_{c}).$ (4)
Typical relaxation functions.
Depending on whether the system falls into Model A or Model B, as defined by Hohenberg and Halperin Hohenberg and Halperin (1977a), and on possible disorder or heterogeneity, the following forms are standard:
  1. 1.
    Exponential: $\phi(t)\!\sim\!e^{-t/\tau}$ (away from $T_{c}$);
  2. 2.
    Stretched exponential (KWW): $\phi(t)\!\sim\!\exp[-(t/\tau)^{\beta_{\mathrm{KWW}}}]$, $0<\beta_{\mathrm{KWW}}<1$ (heterogeneous or glassy systems);
  3. 3.
    Power law: $\phi(t)\!\sim\!t^{-p}$ (at or near $T_{c}$; critical slowing down);
  4. 4.
    Mittag–Leffler (fractional/conformable) Erdelyi (1981): $\phi(t)\!\sim\!E_{\mu}\!\left[-(t/\tau)^{\mu}\right]$, $0<\mu\!\leq\!1$ (broad distributions of time scales).
The one parameter Mittag-Leffler function
$E_{\mu}(-x)=\sum_{k=0}^{\infty}\frac{(-x)^{k}}{\Gamma(1+\mu k)},$ (5)
admits well defined series expansions for small and large arguments. For small arguments $x=(t/\tau)^{\mu}\ll 1$, one obtains
$E_{\mu}\left(-\left[\frac{t}{\tau}\right]^{\mu}\right)=1-\frac{1}{\Gamma(1+\mu)}\left(\frac{t}{\tau}\right)^{\mu}+\frac{1}{\Gamma(1+2\mu)}\left(\frac{t}{\tau}\right)^{2\mu}+O\left(t^{3\mu}\right).$ (6)
Thus, at short times the Mittag Leffler relaxation coincides with the stretched exponential form, because
$E_{\mu}\left(-\left[\frac{t}{\tau}\right]^{\mu}\right)\sim\exp\left[-\frac{1}{\Gamma(1+\mu)}\left(\frac{t}{\tau}\right)^{\mu}\right],\qquad t\ll\tau,$
showing explicitly the link between fractional or conformable relaxation and heterogeneous KWW kinetics.
For large arguments $x\gg 1$, the asymptotic form is
$E_{\mu}(-x)\sim\frac{1}{x\,\Gamma(1-\mu)},\qquad x\to\infty,$ (7)
which corresponds to the slow algebraic decay characteristic of critical slowing down.
These expansions show how a single function interpolates between stretched exponential behavior at short times and a power law decay at long times.
Ergodicity Considerations:
Near critical points, several ergodicity-related issues arise:
  • Ergodicity breaking timescale: The system may appear non-ergodic on experimental timescales as $\tau$ diverges approaching $Tc;$
  • Critical slowing down: The diverging relaxation time means the system takes increasingly long to explore phase space, leading to practical ergodicity breaking;
  • Finite-size effects: In finite systems, true ergodicity is maintained but with timescales $\tau_{erg}$ $\sim L^{z}$ where $L$ is the system size;
[…]
In thermal equilibrium near a continuous transition, the dynamics are (asymptotically) ergodic in finite systems; criticality does not by itself break ergodicity, but the divergent $\tau$ makes mixing arbitrarily slow (“critical slowing down”). Non-ergodicity (or weak ergodicity breaking) arises when additional ingredients are present: quenched disorder/spin-glass physics, constraints, or genuine non-equilibrium protocols (quenches) that produce aging, where two-time correlations $C(t,t_{w})$ lose time-translation invariance. In our conformable setting, $\mu<1$ encodes such slow, broad-spectrum kinetics; equilibrium ergodicity is retained, but relaxation becomes non-Debye and may display aging under driven or quenched conditions. Practically, this means time-averages converge to ensemble averages, yet on time scales that diverge as $\tau\sim|T-T_{c}|^{-z\nu}$.
II Critical scaling and conformable derivatives in physical systems
Near critical points, physical systems exhibit a critical slowing down, where the relaxation time $\tau$ diverges as the correlation length $\xi\sim|T-T_{c}|^{-\nu}$ increases. This behavior is characterized by the dynamic scaling law (Hohenberg and Halperin, 1977b; Goldenfeld, 1992; Amit and Martin-Mayor, 2005)
$\tau\sim\xi^{z}\sim|T_{c}-T|^{-z\nu},$ (8)
where $z$ is the dynamic critical exponent and $\nu$ the correlation length exponent Hohenberg and Halperin (1977b).
Observable thermodynamic quantities exhibit a characteristic power-law behavior in the vicinity of continuous (second-order) phase transitions. Specifically, the heat capacity diverges as
$C(T)\sim|T_{c}-T|^{-\alpha},$ (9)
where $\alpha$ is the critical exponent associated with the specific heat. The spontaneous magnetization follows the relation
$|M(T)|\sim(T_{c}-T)^{\beta},\quad\text{for }T<T_{c},$ (10)
where $\beta$ denotes the critical exponent for the order parameter. Moreover, the magnetic susceptibility diverges according to
$\chi(T)\sim|T_{c}-T|^{-\gamma},$ (11)
where $\gamma$ is the critical exponent characterizing the divergence of the susceptibility. The exponents $\alpha$, $\beta$, and $\gamma$ define the universal scaling behavior near the critical point and are largely independent of the microscopic details of the system. We will show that this scaling naturally introduces temperature-weighted derivatives that match the scaling behavior of critical observables. This property embeds scale invariance directly into the differential structure of the equations, enabling straightforward generation of power-law solutions.
The kinetic coefficient $\Gamma(T)$ plays a central role in the time-dependent relaxation dynamics of systems near criticality. It essentially measures how rapidly a system returns to equilibrium after a small perturbation. Physically, it serves as a mobility-like parameter, often appearing in the Langevin equation or time-dependent Ginzburg–Landau models as a prefactor to the functional derivative of the free energy (Hohenberg and Halperin, 1977b; Kubo, 1966; Forster, 1975; Amit and Martin-Mayor, 2005). Since the kinetic coefficient $\Gamma(T)$ is inversely proportional to the relaxation time, it inherits the corresponding divergence
$\Gamma(T)\sim|T_{c}-T|^{z\nu}.$ (12)
This reflects the critical slowing down of the dynamics near $T_{c}$: the relaxation time diverges, and the kinetic response $\Gamma(T)$ vanishes. In overdamped systems, this corresponds to a loss of mobility, where the system becomes increasingly inert to perturbations near the critical point. Far from $T_{c}$, relaxation is approximately exponential, $\phi(t)\sim e^{-t/\tau}$, but as criticality is approached, $\tau\!\to\!\infty$ and the relaxation crosses over to stretched-exponential Weberszpil (2025a) or power-law forms, characteristic of critical slowing down (Hohenberg and Halperin, 1977b; Goldenfeld, 1992; Godrèche and Luck, 2000; Stanley, 1971). That is, at or near criticality, relaxation is no longer exponential — instead, it becomes power-law, Mittag-Leffer or stretched-exponential (Weberszpil, 2025a) (depending on the model), reflecting critical slowing down and the emergence of long-time tails or non-Markovian behavior.
This work addresses the fundamental question of how conformable derivatives emerge naturally in critical phenomena while offering genuine physical insights beyond classical mean-field theory. We establish rigorous connections between the conformable derivative formalism and established physical mechanisms. First, we demonstrate how anomalous relaxation dynamics near the critical temperature $Tc$ give rise to conformable structures in the adiabatic limit. Second, we show that finite-size effects and boundary constraints lead to temperature-dependent prefactors consistent with the conformable framework. Third, we connect the presence of quenched disorder and spatial inhomogeneities to the emergence of modified scaling laws captured effectively by conformable derivatives. Finally, we relate the conformable exponent to the underlying fractal geometry, interpreting it as an effective dimension governing critical fluctuations.
Remark on the use of critical scaling forms
When expressing the power-law behavior near the critical temperature $T_{c}$, the choice of expression depends on the temperature regime. For temperatures below the critical point ($T<T_{c}$), one uses $(T_{c}-T)^{\beta}$, as this quantity remains positive and correctly describes observables such as the order parameter (e.g., magnetization or superconducting gap), which typically grow as $T$ decreases below $T_{c}$. For temperatures above the critical point ($T>T_{c}$), the appropriate form is $(T-T_{c})^{\beta}$, which applies to quantities that diverge as the system approaches $T_{c}$ from above, such as the heat capacity or magnetic susceptibility. When a symmetric form is desired, valid on both sides of the transition, a common convention is to use the absolute value $|T-T_{c}|^{\beta}$, which appears frequently in generalized scaling laws and in the unified description of critical divergences.
Conformable derivatives: Bridging classical, geometric, and fractional perspectives
As we will show, the conformable derivative provides a powerful intermediary between classical differential operators and fully nonlocal fractional derivatives. While fractional calculus introduces long-range memory and non-locality (Metzler and Klafter, 2000; Evangelista and Lenzi, 2018; Metzler and Klafter, 2004; Mathai and Haubold, 2017), it is often analytically cumbersome and dimensionally ambiguous. The conformable derivative (1) offers a tractable alternative. For $\mu=1$, it reduces to the classical derivative, while for $\mu<1$, it captures essential features of fractional scaling, such as subdiffusion, intermittency, and fractal geometry, without requiring nonlocal integration kernels (Khalil et al., 2014; Abdeljawad, 2015).
This makes the conformable framework especially suitable for systems exhibiting complex thermodynamic behaviors, including those described by generalized statistics (e.g., Tsallis entropy) (Tsallis, 1988; Weberszpil, 2025a, b; Tsallis, 2009; Weberszpil and Helayël-Neto, 2015). Crucially, from a dimensional standpoint, the conformable derivative maintains physical unit consistency,
$D_{T}^{(\mu)}f(T)]=\frac{[f]}{[T]^{\mu}},$ (13)
and all critical exponents derived from it (e.g., $\alpha,\beta$) remain dimensionless. Modified Ginzburg-Landau equations incorporating conformable terms remain compatible with thermodynamic scaling laws and unit balance, confirming both mathematical consistency and physical viability (Weberszpil, 2025b).
Beyond this formal consistency, the conformable derivative admits a compelling geometric interpretation. The operator $D_{T}^{(\mu)}$ effectively acts as a local deformation of the underlying thermodynamic geometry, modifying the metric structure that connects internal energy, entropy, and temperature. In this interpretation, variations in the conformable index $\mu$ alter the curvature of the thermodynamic manifold, introducing a temperature-weighted directional response that encodes local nonequilibrium effects. This approach resonates with the framework of geometrical thermodynamics and the Quantitative Geometrical Thermodynamics (QGT) formalism, in which the state space of thermodynamic variables is endowed with a Riemannian metric reflecting energy–entropy conjugacy and fluctuation geometry (Quevedo and Vázquez, 2008). This quantity can be viewed as a curvature in the space of thermal states, analogous to how spacetime curvature in general relativity modifies geodesics. Here, the thermal evolution is bent by the conformable weighting, making the system’s response scale-sensitive. This captures phenomena such as anomalous diffusion, divergence of correlation length, or emergent power-law scaling near criticality.
Importantly, the framework retains analytical tractability. Unlike renormalization group treatments, conformable models yield closed-form or semi-analytic expressions for key observables such as the order parameter $\psi(T)$, London penetration depth $\lambda_{L}(T)$, and heat capacity $C(T)$, as shown below. These expressions are not only dimensionally coherent but also directly fit experimental data. In particular, fits to superconducting phase transitions recover expected mean-field exponents such as $\beta\approx 1/2$, thereby validating the model’s empirical utility.
In summary, the conformable derivative offers a physically motivated, analytically manageable, and geometrically meaningful framework for modeling critical phenomena. It seamlessly bridges classical dynamics, geometric deformation, and fractional scaling, making it a promising tool for exploring universality, non-locality, and thermodynamic complexity in condensed matter systems and beyond.
To explore how deformed calculus affects the thermodynamic behavior near criticality, we now introduce the conformable derivative framework and show how it modifies standard scaling relations.
III Critical exponents and thermodynamic quantities
In this section, we derive the scaling behavior of key thermodynamic quantities near the critical temperature $T_{c}$ using the conformable derivative framework. This approach yields expressions for the critical exponents $\alpha$, $\beta$, $\gamma$, and $\nu$, which characterize the singular behavior of heat capacity, magnetization, magnetic susceptibility, and correlation length, respectively.
We begin by applying this operator to each observable and derive the corresponding critical exponent in terms of the deformation parameter $\mu$ and characteristic coefficients. These expressions will form the foundation for the quantitative modeling of critical behavior in later sections.
III.1 Phenomenological origin of the conformable critical equation: Heat capacity
To model power-law divergences near critical points, it is common to describe thermodynamic observables such as susceptibility or concentration fluctuations with first-order differential equations, whose solutions yield singularities at the critical temperature $T_{c}$. In the classical framework, such a behavior can be captured by equations of the form (Tinkham, 1996; Landau and Lifshitz, 1978; Goldenfeld, 1992)
$\frac{dC}{dT}=-\frac{\alpha}{T_{c}-T}C(T),$ (14)
which yields the well-known scaling law
$C(T)\sim|T_{c}-T|^{-\alpha}.$ (15)
To incorporate generalized dynamics, such as memory effects, anomalous relaxation, or scaling violations associated with nonextensive systems, we replace the classical derivative with the conformable derivative of order $\mu_{C}\in(0,1]$, defined as
$D_{T}^{(\mu_{C})}f(T):=T^{1-\mu_{C}}\frac{df}{dT}.$ (16)
This local deformation introduces intrinsic scaling into the dynamics while preserving analytic solvability.
The connection between temperature-dependent scaling and dynamics arises from the fact that, near a critical point, the control parameter $T-T_{c}$ governs not only equilibrium singularities but also the time-dependent relaxation of fluctuations. In the framework of dynamic critical phenomena (Hohenberg and Halperin, 1977a; Goldenfeld, 1992), the order-parameter correlation function $\phi(t)$ typically satisfies relaxation-type equations of the form
$\frac{d\phi}{dt}=-\Gamma(T)\,\phi(t),$ (17)
where the kinetic coefficient $\Gamma(T)\!\sim\!(T_{c}-T)^{\nu z}$ vanishes at $T_{c}$, producing critical slowing down. By analogy, the conformable derivative equation
$D_{T}^{(\mu_{C})}C(T)=-\kappa\,(T_{c}-T)^{-1}C(T)$ (18)
introduces an intrinsic scaling between the thermal control parameter $T$ and the effective relaxation dynamics encoded in the fractional exponent $\mu_{C}$. This formal correspondence establishes $T$ as a surrogate evolution variable that parameterizes the deformation of thermodynamic trajectories as the system approaches the critical manifold. Consequently, the conformable formalism embeds the divergence of $C(T)$ within a generalized dynamic framework that preserves analytic solvability while reflecting critical slowing-down behavior.
We thus propose the phenomenological equation for the specific heat,
$D_{T}^{(\mu_{C})}C(T)=-\kappa(T_{c}-T)^{-1}C(T),$ (19)
where $\kappa$ is a dimensionless constant governing the strength of the singularity. Inserting the definition (1) of the conformable derivative, relation (19) becomes
$T^{1-\mu_{C}}\frac{dC}{dT}=-\kappa(T_{c}-T)^{-1}C(T).$ (20)
Division by $C(T)$ on both sides and integration yields
$\int\frac{1}{C(T)}\frac{dC}{dT}\,dT=-\kappa\int T^{\mu_{C}-1}(T_{c}-T)^{-1}\,dT.$ (21)
Near the critical point $T\approx T_{c}$, we make the approximation $T^{\mu_{C}-1}\approx T_{c}^{\mu_{C}-1}$ (valid for $|T-T_{c}|/T_{c}\ll 1$). The integral can then be taken in the sense
$\int\frac{dC}{C}$ $=$ $-\kappa T_{c}^{\mu_{C}-1}\int\frac{dT}{T_{c}-T}$
$\ln C(T)$ $=$ $-\kappa T_{c}^{\mu_{C}-1}\ln|T_{c}-T|+\text{const}$
$C(T)$ $=$ $B|T_{c}-T|^{-\alpha},\quad\alpha=\kappa T_{c}^{\mu_{C}-1}$ (22)
where the integration domain is $T_{0},T]$ with $T_{0}$ sufficiently far from $T_{c}$ such that the approximation holds.
This formulation provides a conformable generalization of classical critical scaling laws, embedding the power-law divergence directly into the modified dynamics. It is particularly suitable for systems exhibiting anomalous thermodynamic behavior, nonlocality, or deviations from classical universality classes.
Collecting the above results, we find the explicit critical form
$C_{V}(T)=B\,|T_{c}-T|^{-\alpha},$ (23)
of the heat capacity near the second-order phase transition, where $B$ is a constant positive amplitude.
The empirical data from physical systems such as niobium often exhibit asymmetric critical behavior, where the divergence of $C_{V}$ is different below and above the critical temperature. This asymmetry arises from distinct microscopic mechanisms in the ordered (superconducting) and disordered (normal) phases, and is supported by both theoretical models and experimental heat capacity curves Stanley (1971). To account for this asymmetry, we adopt the piecewise power-law model
$C_{V}(T)=\begin{cases}B_{1}(T_{c}-T)^{-\alpha_{1}},&T<T_{c}\\[5.69046pt] B_{2}(T_{c}-T)^{-\alpha_{2}},&T>T_{c}\end{cases}$ (24)
for the heat capacity, where $B_{1}$, $B_{2}$, $\alpha_{1}$, and $\alpha_{2}$ are fitting parameters. As we show below, the piecewise expression (24) provides a better fit to real data than a symmetric model and remains analytically tractable within the conformable framework.
Also, for computational purposes, to avoid a divergence exactly at $T=T_{c}$, we introduce a regularization parameter $\epsilon>0$, resulting in the smoothed model
$C_{V}^{\text{reg}}(T)=\begin{cases}B_{1}(|T_{c}-T|+\epsilon)^{-\alpha_{1}},&T<T_{c}\\[5.69046pt] B_{2}(|T_{c}-T|+\epsilon)^{-\alpha_{2}},&T>T_{c}\end{cases}$ (25)
In what follows, the symbol $T_{c}$ denotes the critical temperature associated with the specific phenomenon under consideration. Although its numerical value may differ from one case to another, we retain the same notation $T_{c}$ throughout this work for clarity and to avoid introducing additional symbols.
III.2 Magnetization
We now proceed with the analysis of the magnetization in terms of the deformed operator. We start with the critical form of the magnetization,
$|M|\sim(T_{c}-T)^{\beta}.$ (26)
Application of the concrete form $D_{T}^{\mu_{M}}f(T)=T^{1-\mu_{M}}df(T)/dT$ of the conformable derivative yields
$D_{T}^{(\mu_{M})}|M|=\gamma(T_{c}-T)^{-1}|M|.$ (27)
The resulting differential expression can then be written as
$\frac{d|M|}{|M|}=\gamma T^{\mu_{M}-1}(T_{c}-T)^{-1}\,dT.$ (28)
We integrate both sides. The left-hand side yields
$\int\frac{d|M|}{|M|}=\ln|M|+\text{const.}$ (29)
For the right-hand side, we consider the behavior close to the critical temperature $T\to T_{c}^{-}$. Assuming that $T^{\mu_{M}-1}$ varies slowly as compared to the divergence at $T=T_{c}$, we approximate
$T^{\mu_{M}-1}\approx T_{c}^{\mu_{M}-1}.$ (30)
Hence, the integral becomes
$\int\gamma T^{\mu_{M}-1}(T_{c}-T)^{-1}\,dT$
$\hskip 22.76228pt\approx\gamma T_{c}^{\mu_{M}-1}\int\frac{dT}{T_{c}-T}$
$\hskip 22.76228pt\approx-\gamma T_{c}^{\mu_{M}-1}\ln(T_{c}-T).$ (31)
Substituting back, we obtain
$\ln|M|\approx-\gamma T_{c}^{\mu_{M}-1}\ln(T_{c}-T),$ (32)
witch leads to
$|M|\sim(T_{c}-T)^{\gamma T_{c}^{\mu_{M}-1}}.$ (33)
Identifying the standard critical scaling form $|M|\sim(T_{c}-T)^{\beta}$, we finally obtain the exponent
$\beta=\gamma T_{c}^{\mu_{M}-1}.$ (34)
This expression relates the critical exponent $\beta$ to the temperature-scaling exponent $\mu_{M}$ and the kinetic prefactor $\gamma$, providing a natural thermodynamic link between dynamic behavior and critical ordering.
III.3 Magnetic susceptibility
Also the magnetic susceptibility diverges at criticality,
$\chi(T)\sim|T_{c}-T|^{-\gamma}.$ (35)
This power-law has been treated in both classical and generalized frameworks Tsallis (1988); Plastino and Plastino (1993). Following the same approach as used for the calculation of the heat capacity and magnetization, we write the deformed differential equation for the magnetic susceptibility as
$D_{T}^{(\mu_{\chi})}\chi(T)=-\lambda(T_{c}-T)^{-1}\chi(T).$ (36)
Again, we can conclude that the related exponent $\gamma$ is given by
$\gamma=\lambda T_{c}^{\mu_{\chi}-1}.$ (37)
III.4 Correlation length
The correlation length diverges in the form
$\xi(T)\sim|T_{c}-T|^{-\nu}.$ (38)
Following our approach, we can cast $\xi(T)$ in the form
$D_{T}^{(\mu_{\xi})}\xi(T)=-\rho(T_{c}-T)^{-1}\xi(T),$ (39)
leading to the relation for $\nu$,
$\nu=\rho T_{c}^{\mu_{\xi}-1}.$ (40)
This behavior is central to the renormalization group picture of criticality Goldenfeld (1992) and connects to the scaling approaches pioneered by de Gennes de Gennes (1979) in various critical systems.
III.5 Unified treatment of thermodynamic observables
Following the methodology established for the heat capacity, we apply the conformable framework to other critical observables. Each quantity $Q(T)$ satisfies a deformed differential equation,
$D_{T}^{(\mu_{Q})}Q(T)=-\lambda_{Q}(T_{c}-T)^{-1}Q(T),$ (41)
leading to the critical scaling $Q(T)\sim|T-T_{c}|^{-\epsilon_{Q}}$ with $\epsilon_{Q}=\lambda_{Q}T_{c}^{\mu_{Q}-1}$.
For the magnetic observables, this implies
Magnetization: $\beta=\gamma T_{c}^{\mu_{M}-1}$ (42)
Susceptibility: $\gamma=\lambda T_{c}^{\mu_{\chi}-1}$ (43)
Correlation length: $\nu=\rho T_{c}^{\mu_{\xi}-1}$ (44)
This unified approach eliminates the need for separate phenomenological models while maintaining physical interpretability through the connection between $\mu_{Q}$ and underlying transport or geometric properties.
IV Unified framework and theoretical implications
Building on the conformable formalism, we derive modified evolution equations for thermodynamic observables and extract analytic expressions for the associated critical exponents.
The use of conformable derivatives to model critical phenomena offers the following theoretical and practical advantages.
IV.1 Unified treatment of critical exponents
The model provides a single formalism for expressing all major critical exponents,
$\alpha=\kappa T_{cC}^{\mu_{C}-1},\quad\beta=\gamma T_{cM}^{\mu_{M}-1},$
$\gamma=\lambda T_{c\chi}^{\mu_{\chi}-1},\quad\nu=\rho T_{c\xi}^{\mu_{\xi}-1},$ (45)
where we introduced a distinct notation for the critical temperature related to each physical quantity.
The approach used here avoids the need for distinct phenomenological models for each observable, offering a unified perspective grounded in deformed calculus.
IV.2 Unified thermodynamic framework
The use of deformed or conformable derivatives provides a natural bridge between classical critical phenomena and generalized thermodynamics, particularly within the framework of nonextensive statistics Tsallis (1988); Borges (2004); Weberszpil and Helayël-Neto (2015). By incorporating temperature-dependent scaling, long-range correlations, memory effects, and fractal-like structures, such formulations reproduce anomalous relaxation and power-law distributions that extend beyond the scope of traditional universality classes. The suitability of conformable and $q$-deformed descriptions for systems exhibiting power-law statistics follows from the local scaling properties embedded in their differential operators. When observables obey relations of the type $O(\lambda x)=\lambda^{\eta}O(x)$, the conformable derivative $D^{(\mu)}f(x)=x^{1-\mu}f^{\prime}(x)$ naturally preserves this homogeneity, since its action rescales as $D^{(\mu)}[f(\lambda x)]=\lambda^{\mu}D^{(\mu)}f(\lambda x)$. This property makes it intrinsically compatible with scale-invariant or fractal-like processes, where fractional exponents characterize self-similarity. Furthermore, the conformable operator retains locality in its definition, avoiding the integral memory kernels typical of fractional calculus, while still capturing effective memory and heterogeneity through the scaling index $\mu$. Consequently, once the statistical behavior of a system is known to follow a power law—for example, $p(x)\propto x^{-\alpha}$ or correlation functions $\langle A(0)A(t)\rangle\sim t^{-\gamma}$—the conformable framework provides a compact, analytically tractable description in which the deformation parameter encodes the same scaling exponents that govern the system’s critical or nonextensive dynamics. Importantly, these approaches are not intended to replace the renormalization-group (RG) formalism—which remains the fundamental tool for identifying fixed points and coupling-constant flows—but rather to complement it. Once the underlying system is known to exhibit power-law statistics or scale-invariant fluctuations, conformable and $q$-deformed formulations offer a thermodynamically consistent and analytically transparent representation of the emergent scaling behavior, embedding memory, self-similarity, and heterogeneity into a unified theoretical framework. A major advantage of this approach is its analytic tractability. In contrast to renormalization group techniques, which often require complex iterative procedures, the present model leads to closed-form expressions for critical behavior. The resulting differential equations are easily solvable and lend themselves to direct comparison with experimental data, making the method valuable for both theorists and experimentalists.
Moreover, the conformable or scale-deformed derivatives employed in the formulation naturally introduce power-law solutions. This mirrors the observed scaling of physical observables near the critical temperature $T_{c}$, where critical exponents emerge from the underlying structure of the modified dynamics rather than requiring external assumptions.
Unlike renormalization group methods, this model yields closed-form solutions and is analytically manageable. Its equations can be solved directly or fitted to data without computational overhead, making it practical for both theorists and experimentalists.
To benchmark the conformable model, we now compare its predictions to classical mean-field theory and investigate its behavior in the limiting case where the deformation vanishes.
V Practical advantages over standard approaches
The conformable derivative framework offers several concrete advantages over more conventional scaling analysis and fractional calculus approaches. These are:
Computational Efficiency: Unlike fractional derivatives requiring convolution integrals, conformable derivatives yield algebraic expressions amenable to standard fitting procedures. Computational time scales as $\mathcal{O}(N)$ rather than $\mathcal{O}(N^{2})$ for fractional methods Abdeljawad (2015); Anderson (2016); Khalil et al. (2014). This simplification has made conformable formulations particularly useful in numerical modeling of anomalous diffusion and relaxation, where fractional kernels are computationally costly.
Parameter Interpretability: The deformation parameter $\mu$ directly relates to intrinsic physical properties. For instance, $\mu=1-D_{f}/D$, where $D_{f}$ is the fractal (Hausdorff) dimension and $D$ the embedding dimension, providing a geometric measure of deviation from extensivity and an immediate physical interpretation unavailable in phenomenological scaling fits Weberszpil (2025b); Weberszpil and Helayël-Neto (2015); Anderson (2016). This correspondence links conformable dynamics with fractal metrics and nonextensive thermodynamics, yielding a consistent dimensional interpretation of $\mu$.
Predictive Power: Once calibrated on one observable (e.g., heat capacity), the framework predicts scaling behavior for related quantities (penetration depth, coherence length) without additional fitting. Standard power-law fits treat each observable independently.
Unified Analysis: The approach naturally handles asymmetric scaling above and below $T_{c}$ through different $\mu$ values, capturing microscopic asymmetries that pure mean-field theory cannot address.
VI Modified Ginzburg-Landau equation
Before deriving the Ginzburg–Landau (GL) equation with a conformable derivative and nonlinear interaction term, we first recall the conventional GL-theory. As treated by de Gennes de Gennes (1966), the free energy functional in a $d$-dimensional space reads
$\mathcal{F}[\psi]=\int d^{d}r\left[a(T)|\psi(\mathbf{r})|^{2}+\frac{b}{2}|\psi(\mathbf{r})|^{4}+\frac{1}{2m}\left|\nabla\psi(\mathbf{r})\right|^{2}\right].$ (46)
This classical GL-theory describes the thermodynamic behavior of a complex order parameter $\psi(\mathbf{r})$ near the critical temperature $T_{c}$. The GL free energy functional (46) further involves the temperature-dependent coefficient
$a(T)=a_{0}(T_{c}-T)$ (47)
with $a_{0}>0$, the stabilizing coefficient $b>0$ ensuring the boundedness of the free energy, and the generalized mass (or stiffness parameter) $m$.
To determine the equilibrium configuration, we minimize $\mathcal{F}[\psi]$ with respect to $\psi^{*}(\mathbf{r})$, i.e., $\frac{\delta\mathcal{F}}{\delta\psi^{*}(\mathbf{r})}=0$, which leads us to the Euler–Lagrange equation
$a(T)\psi(\mathbf{r})+b|\psi(\mathbf{r})|^{2}\psi(\mathbf{r})-\frac{1}{2m}\nabla^{2}\psi(\mathbf{r})=0.$ (48)
This nonlinear partial differential equation governs the spatial behavior of the order parameter near criticality. In the case of a spatially uniform system (i.e., $\nabla\psi=0$), relation (48) reduces to the simpler algebraic equation
$a(T)\psi+b|\psi|^{2}\psi=0,$ (49)
whose solutions describe a second-order phase transition and yield the critical exponent $\beta=\tfrac{1}{2}$ under mean-field conditions.
Minimizing the standard GL-free energy functional, it can be shown that the order parameter $\psi(T)$ near the critical temperature $T_{c}$ follows the scaling law
$|\psi(T)|\sim(T_{c}-T)^{\beta},\qquad\text{for }T<T_{c},$ (50)
with the critical exponent $\beta=\tfrac{1}{2}$ under mean-field conditions. By this we mean that we assume that the system is in thermodynamic equilibrium, homogeneous and local (no spatial gradients or long-range correlations), that the fluctuations are negligible, the free energy is analytic near $T_{c}$ and that the order parameter $\psi$ is uniform and real. In the full GL-free energy functional, the gradient term $\frac{1}{2m}|\nabla\psi(\mathbf{r})|^{2}$ is included to account for the energy cost associated with variations of the order parameter $\psi$ in time, temperature, or space. This term is essential in dynamic or spatially inhomogeneous contexts.
In the mean-field approach with vanishing gradient term, $|\nabla\psi(\mathbf{r})|^{2}=0$, the time-independent, spatially uniform GL-free energy density becomes
$\mathcal{F}[\psi]=a(T)|\psi|^{2}+\frac{b}{2}|\psi|^{4},$ (51)
which depends solely on the magnitude of $\psi$ and the temperature $T$. The coefficient $a(T)$ is defined in Eq. (47), and $b>0$. To minimize the free energy, we take the derivative with respect to $\psi^{*}$ (or $\psi$, since it is real-valued) and set it to zero,
$\frac{d\mathcal{F}}{d\psi}=2a(T)\psi+2b|\psi|^{2}\psi=0,$ (52)
such that
$\psi\left[a(T)+b|\psi|^{2}\right]=0.$ (53)
There exit two solutions: (i) $\psi=0$, corresponding to the disordered (symmetric) phase; and (ii) $|\psi|^{2}=-a(T)/b$, which corresponds to the ordered (broken-symmetry) phase for $a(T)<0$, i.e., $T<T_{c}$. Below $T_{c}$, we have $a(T)<0$, so the nontrivial solution reads
$|\psi|^{2}=-\frac{a_{0}(T_{c}-T)}{b}=\frac{a_{0}}{b}(T_{c}-T).$ (54)
Taking the square root, we obtain for positive solution
$|\psi(T)|=\sqrt{\frac{a_{0}}{b}(T_{c}-T)}.$ (55)
Thus, the order parameter vanishes as
$|\psi(T)|\sim(T_{c}-T)^{\beta},\quad\text{with }\beta=\frac{1}{2},$ (56)
$T<T_{C}.$ This is the well known mean-field prediction for the order parameter and is valid under the assumptions stated above. In lower dimensions ($d\leq 3$), fluctuations become important, and this mean-field result typically underestimates the true critical exponent.
VII The conformable GL-model
After establishing the general framework, we apply it to specific thermodynamic quantities—the specific heat and magnetic susceptibility—highlighting the modifications originating from the conformable deformation. We note that in standard GL-theory, spatial variations of the order parameter $\psi(\mathbf{r})$ contribute to the free energy via the gradient term $\frac{1}{2m}|\nabla\psi(\mathbf{r})|^{2}$, which penalizes sharp spatial inhomogeneities and plays a key role in capturing coherence, textures, and defects. This is essential in systems where the spatial coherence or texture of the order parameter matters, such as for vortices in superconductors. In our modified GL framework, however, we are concerned with the thermal evolution of the order parameter $\psi(T)$, i.e., with the description of how the this order parameter evolves as a function of temperature $T$, especially in the vicinity of the critical point (critical temperature) $T_{c}$. To this end we assume that the system is homogeneous in space.
To build a thermodynamic analog of the GL-theory, we formally treat the temperature $T$ as the independent coordinate, akin to space $\mathbf{r}$ in the conventional setting. In this case, we introduce a conformable (deformed) derivative with respect to temperature, denoted by
$T_{\alpha}\psi(T):=T^{1-\alpha}\frac{d\psi}{dT},$ (57)
which naturally embeds a temperature-dependent scaling. The conformable operator acts as a thermal gradient and captures the power-law behavior typical of critical phenomena. The conformable operator (57) generalizes the usual derivative and introduces a natural temperature-dependent scaling, allowing the model to capture anomalous thermodynamic behavior while remaining analytically tractable. The corresponding modified GL-free energy then becomes
$\mathcal{F}_{\alpha}[\psi]=\int\left[a(T)|\psi|^{2}+\frac{b}{2}|\psi|^{4}+\frac{1}{2m}|T_{\alpha}\psi(T)|^{2}\right]dT,$ (58)
where $a(T)$ has the same form (47) to encode the proximity to the critical point, $b>0$ ensures stability, and the last term mimics the kinetic contribution, now expressed in thermal rather than spatial form. The parameter $m$ retains its interpretation as a stiffness constant, and $\alpha\in(0,1]$ quantifies the deviation from classical scaling. For $\alpha=1$, the standard GL-theory is recovered.
The formulation (58) of the free energy functional is central to this work and preserves the physical dimensions of energy density, and it ensures that all critical exponents derived from it are dimensionless and compatible with thermodynamic scaling theory. Furthermore, the temperature-dependent weighting introduced by the factor $T^{1-\alpha}$ can be interpreted geometrically as a curvature in thermal space, analogous to how curved metrics affect geodesics in general relativity. It effectively embeds power-law sensitivity into the response functions of the system.
To minimize the free energy, we use the variational principle $\delta\mathcal{F}_{\alpha}/\delta\psi^{*}=0$, which yields the modified Ginzburg–Landau equation:
$a(T)\psi+b|\psi|^{2}\psi-\frac{1}{2m}T^{2(1-\alpha)}\frac{d^{2}\psi}{dT^{2}}-\frac{(1-\alpha)}{m}T^{1-2\alpha}\frac{d\psi}{dT}=0.$ (59)
Here $a(T)=a_{0}(T/T_{c}-1)$ retains the usual temperature dependence, while the conformable terms introduce explicit temperature-weighted derivatives. The order parameter $\psi$ describes the macroscopic condensate amplitude, and near the critical temperature it follows the scaling law
$\psi(T)\simeq\psi_{0}\left(1-\frac{T}{T_{c}}\right)^{\beta},$ (60)
where $\psi_{0}$ is the equilibrium amplitude at $T=0$ and $\beta$ is the critical exponent associated with the order-parameter vanishing as $T\to T_{c}$. The dependence $\beta=\beta(\alpha)$ reflects how the conformable index $\alpha$ modifies the effective critical behavior of the system. This expression recovers the expected behavior from mean-field theory when $\alpha=1$, but allows deviations due to the deformation parameter. The above equations can also be obtained via a variational formulation adapted to deformed derivatives, as discussed in Weberszpil and Helayël-Neto (2016).
The nonlinear term $|\psi|^{2}\psi$ contributes to the saturation of the order parameter as $T\to 0$, while the conformable kinetic term introduces corrections near $T_{c}$, mimicking effects from spatial or temporal fluctuations. The formulation presented here enables capturing intermediate critical dynamics and non-locality without invoking full fractional calculus or spatial disorder Tinkham (2004); Bardeen et al. (1957); Gorter and Casimir (1934); Park and Dorsey (2003), making it suitable for describing second-order phase transitions such as superconductivity. In this sense, the conformable GL-model complements traditional models such as BCS theory Bardeen et al. (1957), the two-fluid model Gorter and Casimir (1934), and the phenomenological framework detailed in Tinkham (2004). For systems with strong fluctuations and short coherence length, additional comparison with the work in Park and Dorsey (2003) provides insight into limitations and refinements.
A key quantity in critical phenomena is the correlation length, whose divergence governs universality. We now derive its scaling law within the conformable approach and discuss physical implications.
VIII Application to superconducting phase transitions
Superconducting phase transitions, especially in type-II and high-$T_{c}$ materials, are characterized by critical phenomena that can be reformulated using conformable derivatives. This includes the temperature behavior of the order parameter, coherence length, penetration depth, and heat capacity.
VIII.1 Simplified dynamics near the critical point: Connection to the conformable derivative framework
We begin with the full conformable GL-equation (59) obtained from minimizing the deformed free energy functional. This nonlinear second-order differential equation governs the thermal evolution of the order parameter $\psi(T)$, incorporating both nonlinear saturation ($|\psi|^{2}\psi$) and conformable kinetic effects through temperature-dependent coefficients.
Near the critical point, the system undergoes a continuous phase transition. In this regime, we make the following approximations to simplify the analysis: (i) We assume that $\psi(T)$ is real-valued and slowly varying. This allows us to treat $\psi$ and its derivatives in a simplified manner. Moreover, we (ii) neglect higher-order kinetic corrections. The second-order derivative term and the prefactor corrections from conformable calculus are of higher order and can be dropped close to $T_{c}$. This reduces the second-order equation to a first-order approximation. Finally, we (iii) retain only the dominant balance between the kinetic and the potential terms. Taking into account these approximations, the reduced equation can be cast as
$\frac{d\psi}{dT}=a(T)\psi-b|\psi|^{2}\psi,$ (61)
where $a(T)$ is defined in Eq. (47) and $b>0$ ensures the stability of the ordered phase.
The modified Ginzburg–Landau equation introduced here describes the temperature-dependent behavior of the order parameter near the critical point. The term “time-dependent” has been replaced by “temperature-dependent” to emphasize that the conformable
operator acts on the thermal variable $T$, rather than on real time $t$.
VIII.2 Emergence of critical scaling from conformable GL relaxation with temperature-dependent kinetics
While the standard Ginzburg-Landau (GL) equation near the critical temperature $T_{c}$ yields a regular (non-singular) behavior for the order parameter $\psi(T)$, a phenomenological equation capable of reproducing power-law critical behavior can be obtained by introducing a thermally deformed kinetic coefficient into the GL-like relaxation framework. This modification effectively incorporates the temperature-dependent slowing down of dynamics near $T_{c}$ into the evolution of the order parameter.
As before, we generalize the time derivative to the form
$\frac{d\psi}{dT}\to D_{T}^{(\mu_{\psi})}\psi(T):=T^{1-\mu_{\psi}}\frac{d\psi}{dT},$ (62)
which reflects the modified temperature response of the system in the conformable thermodynamic formulation.
At the heart of the conformable dynamics approach lies a generalized time-dependent GL-type equation that governs the thermal evolution of the order parameter $\psi(T)$ under non-equilibrium conditions. This equation introduces a thermally modulated kinetic response and is given by
$\Gamma(T)\,T^{1-\mu_{\psi}}\frac{d\psi}{dT}$ $=-\frac{\delta\mathcal{F}}{\delta\psi},$ (63)
where $\Gamma(T)$ is a temperature-dependent kinetic coefficient encoding the critical slowing down, $\mu_{\psi}$ is the conformable deformation parameter, and $\mathcal{F}[\psi]$ is the free energy functional. Relation (63) is not merely a phenomenological extension but constitutes the cornerstone of the proposed non-equilibrium conformable dynamics framework. It encapsulates the scale-sensitive kinetic behavior characteristic of critical systems, embedding a power-law response directly into the governing evolution law. Equation (63) allows the derivation of critical exponents such as $\beta$ from first principles within a thermodynamically consistent and geometrically interpretable formalism, thereby bridging equilibrium mean-field results with more general, deformation-induced scaling behaviors.
We start with the standard GL-potential (neglecting gradient terms)
$\mathcal{F}[\psi]=\int\left[a_{0}(T_{c}-T)|\psi|^{2}+\frac{b}{2}|\psi|^{4}\right]dT,$ (64)
which produces the variational derivative
$\frac{\delta\mathcal{F}}{\delta\psi}=a_{0}(T_{c}-T)\psi+b|\psi|^{2}\psi.$ (65)
Substituting this expression into Eq. 63 yields
$\Gamma(T)T^{1-\mu_{\psi}}\frac{d\psi}{dT}=-a_{0}(T_{c}-T)\psi-b|\psi|^{2}\psi.$ (66)
Near the critical point, $T\to T_{c}$, the order parameter $\psi(T)$ becomes small, allowing us to neglect the nonlinear term in Eq. (66). The relaxation equation then reduces to the linearized form
$\Gamma(T)\,T^{1-\mu_{\psi}}\frac{d\psi}{dT}\approx-a_{0}(T_{c}-T)\psi.$ (67)
Rewriting this in differential form yields the form
$\frac{d\psi}{\psi}=-\frac{a_{0}}{\Gamma(T)}T^{\mu_{\psi}-1}(T_{c}-T)\,dT,$ (68)
which will be used to extract the scaling behavior of the order parameter close to the critical temperature.
To obtain the desired critical scaling behavior, we now propose that the kinetic coefficient $\Gamma(T)$ vanishes near $T_{c}$, reflecting the critical slowing down. Specifically, we take
$\Gamma(T)=\frac{1}{\gamma}(T_{c}-T)^{2},$ (69)
where $\mu_{\psi}\in(0,1]$ is a conformable deformation parameter, and where $\gamma$ is a dimensional constant. Substituting back, we obtain the phenomenological differential equation
$\frac{d\psi}{\psi}=-\gamma T^{\mu_{\psi}-1}(T_{c}-T)^{-1}\,dT.$ (70)
This equation directly yields the critical behavior
$\int\frac{d\psi}{\psi}=-\gamma T^{\mu_{\psi}-1}\int\frac{dT}{(T_{c}-T)}.$ (71)
Integration produces the result
$\ln\psi(T)=\gamma T_{c}^{\mu-1}\ln|T-T_{c}|+\text{const.}$ (72)
From this we conclude that
$\psi(T)\sim(T_{c}-T)^{\beta(\mu)},\quad\text{with}\quad\beta(\mu)=\gamma T_{c}^{\mu_{\psi}-1},$ (73)
where we approximated $T^{\mu_{\psi}-1}\approx T_{c}^{\mu_{\psi}-1}$ near $T=T_{c}$. Thus, the critical exponent $\beta$ emerges naturally from a GL-type relaxation model with thermally modulated kinetics.
Here, the scaling exponent $\beta=\tfrac{1}{2}$ emerges in the equilibrium case (in the adiabatic limit $\frac{d\psi}{dT}\to 0$ or at least for the stationarity condition dominated by the free energy balance as detailed in Sec. VI) and App. B), corresponding to the local, undeformed limit ($\alpha=1$). The case $\beta>\tfrac{1}{2}$ arises when the conformable deformation ($\mu=2\alpha<1$) introduces a generalized thermodynamic response. The general form for the critical exponent $\beta(\mu)$ can be viewed as a function encoding microscopic memory or a fractal behavior in temperature evolution.
Deviations from the mean-field behavior occur when (i) thermal fluctuations near $T_{c}$ become significant (especially in $d<4$), requiring renormalization group treatment; (ii) nonlocal effects or memory are present, as modeled here via conformable derivatives; (iii) the system is driven out of equilibrium or experiences dissipation or long-time correlations. These effects modify the scaling balance in the GL-equation and yield $\beta>\tfrac{1}{2}$, consistent with experimental observations in real superconductors—which often exhibit a critical behavior closer to the 3D XY universality class ($\beta\approx 0.35$) (Campostrini et al., 2001; Goldenfeld, 1992; Hohenberg and Halperin, 1977a; Osborn et al., 2003).
VIII.3 Emergence of penetration-depth scaling exponent from thermally deformed conformable dynamics
The London penetration depth $\lambda_{L}(T)$ characterizes the distance over which magnetic fields decay inside a superconductor. Near the critical temperature $T_{c}$, $\lambda_{L}(T)$ diverges as superconductivity breaks down. In analogy with critical phenomena, this divergence can be modeled by a deformed dynamical relation using conformable derivatives.
We begin with the assumption that $\lambda_{L}(T)$ obeys a conformable differential equation of the form
$D_{T}^{(\mu)}\lambda_{L}(T)=-\gamma\,(T_{c}-T)^{-1}\lambda_{L}(T),$ (74)
where $\mu\in(0,1]$ is the order of the deformation and $\gamma$ is a positive scaling constant. Equation (74) states that the relative temperature derivative of $\lambda_{L}$, modulated by $T^{1-\mu}$, is proportional to the inverse distance from the critical point $T_{c}$. It encapsulates both the scaling nature of the divergence and possible memory/fractal effects through $\mu$.
We proceed by first substituting the definition of the conformable derivative,
$T^{1-\mu}\frac{d\lambda_{L}}{dT}=-\gamma(T_{c}-T)^{-1}\lambda_{L}(T)$ (75)
such that
$\frac{d\lambda_{L}}{\lambda_{L}}=-\gamma\,T^{\mu-1}(T_{c}-T)^{-1}dT.$ (76)
Near the critical point $T\approx T_{c}$, we approximate $T^{\mu-1}\approx T_{c}^{\mu-1}$, yielding
$\frac{d\lambda_{L}}{\lambda_{L}}\approx-\gamma\,T_{c}^{\mu-1}(T_{c}-T)^{-1}dT.$ (77)
Integration on both sides produces
$\ln\lambda_{L}(T)=-\gamma T_{c}^{\mu-1}\ln|T-T_{c}|+\text{const.}$ (78)
Thus we obtain the scaling law
$\lambda_{L}(T)=B\,|T-T_{c}|^{-\alpha},\quad\text{with}\quad\alpha=\gamma T_{c}^{\mu-1}.$ (79)
This result shows that the penetration depth diverges as a power-law near $T_{c}$, with an exponent that depends on the conformable order $\mu$, the critical temperature $T_{c}$, and the coupling constant $\gamma$.
To include a possible asymmetry in the behavior above and below $T_{c}$, we generalize the penetration depth to a piecewise form with different exponents, in a analogy to our procedure with $C_{V}$ above,
$\lambda_{L}(T)=\begin{cases}B_{1}\,|T-T_{c}|^{-\alpha_{1}},&T<T_{c},\\[4.0pt] B_{2}\,|T-T_{c}|^{-\alpha_{2}},&T\geq T_{c},\end{cases}$ (80)
where $\alpha_{i}=\gamma_{i}T_{c}^{\mu_{i}-1}$.
Finally, to prevent a divergence and ensure continuity near $T=T_{c}$, we introduce a regularization parameter $\epsilon>0$, resulting in the practical model used in the fit procedure below,
$\lambda_{L}(T)=B_{i}\left(|T-T_{c}|+\epsilon\right)^{-\alpha_{i}},$ (81)
for $i=1,2$ above and below $T_{c}$, see Eq. (80). This regularized conformable model reproduces the observed smooth divergence and captures the asymmetric critical behavior seen in (finite) superconducting materials such as niobium.
Connection between penetration depth and order parameter
The London penetration depth $\lambda_{L}(T)$ characterizes the distance over which an external magnetic field decays exponentially inside a superconductor. It provides a direct measure of the material’s ability to expel a magnetic flux via the Meissner effect. In GL-theory, the superconducting state is described by the complex order parameter $\psi(T)$, whose squared magnitude $|\psi(T)|^{2}$ corresponds to the density of the superconducting carriers. This connection leads to a fundamental relation between the order parameter and the electromagnetic response of the superconductor (Tinkham, 1996; De Gennes, 1999).
From the London equation, which relates the supercurrent $\mathbf{J}_{s}$ to the vector potential $\mathbf{A}$, we have (London and London, 1935; Tinkham, 1996)
$\nabla\times\mathbf{J}_{s}=-\frac{n_{s}e^{2}}{m}\mathbf{B},$ (82)
where $n_{s}$ is the superconducting carrier density, $e$ is the elementary charge, $m$ is the effective mass, and $\mathbf{B}$ is the magnetic field. The London penetration depth is then given by the expression
$\lambda_{L}^{-2}(T)=\frac{\mu_{0}n_{s}(T)e^{2}}{m},$ (83)
where $\mu_{0}$ is the vacuum permeability.
In the GL-framework, the carrier density is proportional to the square of the order parameter, $n_{s}(T)\sim|\psi(T)|^{2}$. Substituting into the expression for $\lambda_{L}^{-2}(T)$, we obtain the key relation
$\lambda_{L}^{-2}(T)\propto|\psi(T)|^{2}.$ (84)
This equation shows that, as the temperature approaches its critical value $T_{c}$, the order parameter $\psi(T)\to 0$, and consequently $\lambda_{L}(T)\to\infty$. Physically, this reflects the loss of superconductivity and the full penetration of magnetic fields into the material. Hence, the temperature dependence of $\lambda_{L}(T)$ provides a direct experimental probe of the order parameter’s behavior and is commonly used to extract critical exponents and validate theoretical models.
Near the critical temperature $T_{c}$, we have shown from the conformable GL-equation that the superconducting order parameter behaves as
$|\psi(T)|\sim\left(1-\frac{T}{T_{c}}\right)^{\beta},\qquad\text{for }T<T_{c},$ (85)
where the critical exponent $\beta$ depends on the conformable deformation parameter $\alpha$. This reduces to the standard mean-field behavior with $\beta=\tfrac{1}{2}$ when $\alpha=1$, recovering the classical GL-result Tinkham (1996). Thus, near the critical temperature $T_{c}$, the superconducting order parameter vanishes in power-law form. Moreover, the London penetration length according to relation (84) scales as
$\lambda_{L}(T)\sim(T_{c}-T)^{-\beta}.$ (86)
Thus the penetration depth diverges as the temperature approaches $T_{c}$ from below. This scaling relation is a direct consequence of the GL-formalism and can be used to extract the critical exponent $\beta$ from experimental measurements of $\lambda_{L}(T)$ near the superconducting transition.
VIII.4 Conformable scaling of the specific heat near $T_{c}$: Heat capacity jump
In classical thermodynamics of phase transitions, the specific heat near the critical temperature $T_{c}$ typically scales as (Stanley, 1971; Goldenfeld, 1992)
$C(T)\sim|T-T_{c}|^{-\alpha},$ (87)
where $\alpha$ is the associated critical exponent. In mean-field theory, $\alpha=0$, corresponding to a finite jump at $T_{c}$ (Landau and Lifshitz, 1978; De Gennes, 1999). However, in more generalized or fluctuation-driven models, $\alpha$ may take nonzero values and describe a divergence (Fisher, 1967; Hohenberg and Halperin, 1977a; Campostrini et al., 2001).
To generalize the scaling behavior (87) within the conformable derivative framework, we showed that the specific heat satisfies a deformed differential equation of the form of Eq. (19), such that we find the scaling form
$C(T)\sim|T-T_{c}|^{-\alpha},\qquad\text{with}\qquad\alpha=\kappa T_{c}^{\mu_{C}-1}.$ (88)
Thus the critical exponent $\alpha$ is no longer a universal constant but depends explicitly on the deformation parameter $\mu_{C}$ and the critical temperature $T_{c}$. The conformable framework thus naturally introduces a generalized scaling law for the specific heat.
The generalized conformable framework enables direct fitting of experimental data from superconducting systems using critical scaling laws derived from deformed thermodynamic equations, as shown in the next Section. Specifically, the temperature dependence of key observables can be analyzed as follows: the order parameter $\psi(T)$ may be extracted from angle-resolved photoemission spectroscopy (ARPES), tunneling spectroscopy, or other coherence-sensitive measurements; the London penetration depth $\lambda_{L}(T)$ is accessible through microwave cavity perturbation, magnetic susceptibility, or muon spin rotation techniques; and the specific heat $C(T)$ is measured via calorimetry near the superconducting transition. Each observable exhibits a specific power-law scaling near the critical temperature $T_{c}$, governed by exponents that emerge naturally from the conformable derivative formulation.
Summarizing these results, we note that when solving the time-dependent or inhomogeneous problem in the context of GL-theory, derivative terms must be kept and may dominate if $\psi(T)$ varies sharply. However, for static, near-equilibrium (mean-field) behavior, they contribute only to subleading corrections. The Landau-based equation is an equilibrium condition: it comes from minimizing a free energy, assuming time-independence, homogeneity, and absence of fluctuations. The conformable dynamic equation, instead, describes a non-equilibrium evolution, in which the system may be relaxing toward equilibrium with scale-sensitive kinetics due to $\mu_{\psi}\neq 1$. Therefore, the deformed equation gives rise to a modified critical exponent, which reduces to $\beta=1/2$ only when $\mu_{\psi}=1$, i.e., when the dynamics are classical. The exponent $\beta=1/2$ emerges from equilibrium minimization of the GL-free energy, while $\beta=\gamma T_{C}^{\mu_{\psi}-1}$ results from integrating the conformable dynamic evolution equation, which accounts for non-equilibrium scaling governed by the conformable derivative. They agree only when the deformation vanishes.
Figure 1: Heat capacity of niobium fitted by the smoothed conformable piecewise model (25) with $C_{V}(T)=B_{i}(|T-T_{c}|+\epsilon)^{-\alpha_{i}}$. Fitted parameters with 95% confidence intervals: $B_{1}=0.411\pm 0.023$, $\alpha_{1}=0.644\pm 0.031$ ($T<T_{c}$); and $B_{2}=0.808\pm 0.041$, $\alpha_{2}=0.567\pm 0.028$ ($T>T_{c}$). We used the smoothing factor $\epsilon=0.293\pm 0.015$ to ensure regularization near the divergence at the critical temperature $T_{c}=8.700\pm 0.005$ K. Data from Brown et al. (1953). The fitting employed weighted least-squares with weights proportional to $1/\sigma_{i}^{2}$ where $\sigma_{i}$ are experimental uncertainties.
IX Numerical fitting and data analysis
To test the predictive power of the conformable formalism, we apply it to experimental data for the superconducting transition of niobium.
IX.1 Heat capacity of niobium superconductor
Figure 1 shows the experimental heat capacity data for niobium near its superconducting transition, along with a conformable piecewise fit that captures the critical behavior on both sides of the transition temperature $T_{c}=8.7$ K. The fitted parameters are provided in the caption. For the fit we used the regularized model (25). While the fit is not perfect, it demonstrates good agreement with the data over a wider interval around the critical temperature.
IX.2 Penetration depth and conformable scaling
The temperature dependence of the London penetration depth $\lambda_{L}(T)$ provides key informations about the superconducting coherence and the fluctuation behavior near the critical temperature $T_{c}$. Analogous to the divergence observed in the specific heat, the penetration depth exhibits a critical-like behavior and can be modeled using the conformable scaling framework.
We adopt the piecewise model (80) with the regularized form (81), where $\epsilon>0$ is the smoothing parameter that ensures a regularized behavior at the transition point. The conformable derivative framework again justifies this scaling law via the modified differential relation (74) such that $\lambda_{L}(T)\sim|T-T_{c}|^{-\gamma T_{c}^{\mu-1}}$.
Figure 2: Fit of the normalized London penetration depth $\lambda_{L}(T)/\lambda_{0}$, where $\lambda_{0}$ denotes the London penetration depth extrapolated to zero temperature, serving as a normalization constant for comparison between theory and experiment. The data for niobium are fitted using the conformable piecewise model [Eq. (80)] with the regularized form $\lambda_{L}(T)=B_{i}\left(|T-T_{c}|+\epsilon\right)^{-\alpha_{i}},$ as given in Eq. (81). The fitted parameters are $B_{1}=1.57$, $\alpha_{1}=0.68$ for $T<T_{c}$ and $B_{2}=5.01$, $\alpha_{2}=0.98$ for $T\geq T_{c}$, with $\epsilon=0.055$ and $T_{c}=9.78\penalty 10000\ \text{K}$. Experimental data are taken from Maxfield and McLean (1965).
The temperature dependence of the normalized London penetration depth for niobium is shown in Fig. 2. The data are well described by the conformable piecewise model with distinct scaling exponents below and above the critical temperature. The fitted parameters are listed in the caption of Fig. 2.
IX.3 Coherence length from $H_{c2}(T)$ in niobium
To further validate our theoretical predictions, we turn to experimental data on niobium and extract the temperature-dependent coherence length using the upper critical field $H_{c2}(T)$.
IX.3.1 Theoretical background
In the GL-framework, the upper critical magnetic field $H_{c2}(T)$ is related to the superconducting coherence length $\xi(T)$ by the expression Werthamer et al. (1966); Tinkham (1996); De Gennes (1999)
$H_{c2}(T)=\frac{\Phi_{0}}{2\pi\xi^{2}(T)},$ (89)
where $\Phi_{0}=h/2e\approx 2.07\times 10^{-15}$ Wb is the magnetic flux quantum. This inverse-square relation permits the extraction of $\xi(T)$ from experimental measurements of $H_{c2}(T)$.
Near the critical temperature $T_{c}$, the coherence length diverges according to a power law [Eq. (38)]. Such scaling behavior arises naturally when the Ginzburg–Landau framework is extended to include fractional or conformable spatial derivatives. These generalized operators account for nonlocality, fractality, and memory in the order parameter field, yielding a power-law divergence of the correlation length and critical exponents beyond mean-field values Laskin (2000); Tarasov (2010); Weberszpil and Helayël-Neto (2015, 2016); West (2020). Similar approaches using fractional or stochastic generalizations of the GL-equation have also been explored in dynamical settings Zhang et al. (2022). To see this, consider the modified GL-free energy with the conformable derivative
$|D_{x}^{(\mu)}\psi(x)|^{2}=x^{2(1-\mu)}\left(\frac{d\psi}{dx}\right)^{2},$ (90)
which replaces the standard gradient term $|\nabla\psi|^{2}$. The corresponding GL-equation in one dimension becomes
$a_{0}(T-T_{c})\psi+b|\psi|^{2}\psi$
$-\frac{1}{2m}\left[x^{2(1-\mu)}\frac{d^{2}\psi}{dx^{2}}+(1-\mu)x^{1-2\mu}\frac{d\psi}{dx}\right]=0.$ (91)
Assuming a linearized form near $T_{c}$ and a trial solution $\psi(x)=\psi_{0}e^{-x/\xi}$, we compute the derivatives
$\frac{d\psi}{dx}$ $=-\frac{\psi_{0}}{\xi}e^{-x/\xi},$ (92)
$\frac{d^{2}\psi}{dx^{2}}$ $=\frac{\psi_{0}}{\xi^{2}}e^{-x/\xi}.$ (93)
Substituting these into the kinetic term (90) produces
$\text{Kinetic term}\approx\frac{1}{2m}\left[\frac{x^{2(1-\mu)}}{\xi^{2}}+\frac{(1-\mu)x^{1-2\mu}}{\xi}\right]\psi(x).$ (94)
Balancing with the linear term $a_{0}(T-T_{c})\psi$ in relation (91) we find
$a_{0}(T-T_{c})\sim\frac{1}{2m}\left[\frac{x^{2(1-\mu)}}{\xi^{2}}+\frac{(1-\mu)x^{1-2\mu}}{\xi}\right].$ (95)
Let $\tau=T_{c}-T>0$ be the reduced temperature. The modified Ginzburg–Landau (GL) equation involves two characteristic contributions: the standard term $T_{1}(\tau,\xi)\propto\xi^{-2}$, associated with the spatial gradient energy, and the conformable (or fractional) correction $T_{2}(\tau,\xi)\propto\tau$, which encodes temperature-dependent scaling.
The relative magnitude of these two terms determines the dominant physical regime. The crossover point is obtained when both contributions are of comparable strength,
$T_{1}(\tau_{*},\xi_{*})=T_{2}(\tau_{*},\xi_{*}),$ (96)
which defines a characteristic temperature scale $\tau_{*}$ and a corresponding coherence length $\xi_{*}$. In practice, $\tau_{*}$ marks the boundary between the mean-field region and the conformable-dominated region, and $\xi_{*}$ is the coherence length at this crossover.
As $T\to T_{c}$, the reduced temperature becomes very small ($\tau\ll\tau_{*}$) and the coherence length grows ($\xi\gg\xi_{*}$). In this regime, the conformable correction dominates, leading to the scaling law
$\xi(T)\sim(T_{c}-T)^{-1}.$ (97)
In the opposite regime, defined by $\tau\gg\tau_{*}$ (equivalently $\xi\ll\xi_{*}$), the standard GL term prevails, and the classical mean-field result is recovered,
$\xi(T)\sim(T_{c}-T)^{-1/2}.$ (98)
Therefore, $(\tau_{*},\xi_{*})$ act as crossover parameters separating two asymptotic domains: a critical region governed by conformable scaling and a mean-field region governed by conventional GL behavior. Thus, the power-law divergence (38) is recovered with an effective exponent $\nu$ that may vary depending on which term dominates and the value of the conformable parameter $\mu$. In fractional generalizations of the GL-theory, one may therefore consider replacing the usual spatial gradient term $|\nabla\psi|^{2}$ by this deformed expression involving the conformable derivative, leading to a modified scaling law for $\xi(T)$.
The experimental data for $H_{c2}(T)$ were extracted from Fig. 4 of the classical work by Williamson Williamson (1970), which presents the behavior of the upper critical field for niobium. The experimental points (open circles) were digitized using the WebPlotDigitizer tool.
The original figure displays the data in a dimensionless form, with the vertical axis representing the ratio $H_{c2}^{2}/\left(dH_{c2}/dt\right)_{t=1}$ as a function of the reduced temperature $t=T/T_{c}$. To reconstruct the actual values of $H_{c2}(T)$ in physical units (Tesla), we assumed a functional dependence near the critical transition of the form
$H_{c2}(T)=H_{c2}(0)\left(1-\frac{T}{T_{c}}\right)^{n},$ (99)
where $H_{c2}(0)$ denotes the extrapolated upper critical field at zero temperature and $n$ is a fitting exponent. The critical temperature was fixed at $T_{c}=9.2$ K, consistent with values for pure niobium and in line with the parameters adopted in literature Tinkham (1996). After conversion, the data were fitted with the empirical model (99), enabling the subsequent extraction of the temperature dependence of the coherence length $\xi(T)$ through inversion of relation (89), i.e.,
$\xi(T)=\sqrt{\frac{\Phi_{0}}{2\pi H_{c2}(T)}}.$ (100)
This procedure provides an experimental foundation to assess the validity of theoretical models for the temperature dependence of the coherence length.
We note that here we claim that the use of a conformable derivative in the GL-framework is motivated by the need to capture mesoscopic and spatially inhomogeneous effects that are not adequately described by the standard theory. In particular, real superconducting samples often exhibit intrinsic disorder, nonuniform pinning landscapes, or microstructural granularity, leading to effective nonlocal interactions and spatially varying coherence. The conformable derivative introduces a scale-dependent deformation of the spatial gradient, which can mimic the impact of fractal-like geometries or finite-size scaling in constrained domains. This modification allows for a more flexible description of the kinetic term and leads to a generalized coherence length scaling, making the theory more compatible with experimental deviations from classical mean-field behavior.
IX.3.2 Results and analysis
Figure 3: Extracted coherence length $\xi(T)$ (blue dots) and fitted model (red line) using $\xi(T)=A(T_{c}-T)^{-\nu}$ with $\nu=0.329$ and $A=40.76$. Experimental data adapted from Williamson (1970), see text.
Figure 3 displays the digitized data and the fitted model. The excellent agreement indicates the suitability of the power-law model in this temperature range. The coherence length $\xi(T)$ describes the spatial extent over which the superconducting order parameter $\psi$ varies appreciably. As $T\to T_{c}$, $\xi(T)\to\infty$, and $A$ sets the overall magnitude of $\xi$ for temperatures below $T_{c}$. This coefficient $A$ encodes material-specific properties, including microscopic interactions, effective mass of Cooper pairs, and strength of the pairing potential.
The extracted critical exponent $\nu\approx 0.329$ is notably smaller than the classical GL-mean-field prediction of $\nu=1/2$. Several factors may contribute to this deviation: (i) A limited resolution or experimental uncertainty in the high-temperature region ($T\to T_{c}$); (ii) residual disorder or impurity-induced scattering effects not incorporated into the ideal GL-framework; (iii) memory effects—either thermal or spatial—captured through conformable (fractional-order) derivatives; (iv) effective averaging over anisotropic coherence lengths in polycrystalline or textured samples.
From the expression
$\xi(T)=A(T_{c}-T)^{-\nu},$ (101)
we deduce the units of the coefficient $A$ by dimensional analysis. Since $\xi(T)$ is given in nanometers (nm), $(T_{c}-T)$ in Kelvin (K), and $\nu$ is dimensionless, $A$ must have units of $\text{nm}\cdot\text{K}^{\nu}$. Given the fitted exponent $\nu\approx 0.329$ and $A=40.76$, this means that
$\xi(T)\approx 40.76(T_{c}-T)^{-0.329}\quad\text{[in nm]}.$ (102)
To express $A$ in SI units (meters), we convert according to
$A_{\text{SI}}=40.76\times 10^{-9}\ \text{m}\cdot\text{K}^{0.329}=4.076\times 10^{-8}\ \text{m}\cdot\text{K}^{0.329}.$ (103)
Such a reduced exponent is qualitatively compatible with generalized GL-models that incorporate non-integer-order differential structures.
For clean elemental superconductors like niobium, coherence lengths on the order of 30-40 nm are typical at intermediate temperatures. The value $A=40.76nm\cdot\text{K}^{\nu}$ is thus physically reasonable and consistent with known properties of niobium. It represents the critical amplitude of the coherence length near the critical temperature $T_{c}$. Its physical meaning and units depend on how the data was scaled in the fitting process. A larger $A$ implies that $\xi(T)$ becomes large faster as $T\to T_{c}$, for a fixed $\nu$. This can be influenced by material purity, anisotropy or also electron-phonon coupling strength.
X Conclusions and outlook
In this work, we developed and applied the conformable derivative framework to describe critical phenomena near continuous phase transitions. By deforming the standard derivative operator through a temperature- or space-dependent scaling, we derived modified evolution equations for key thermodynamic observables, including the heat capacity, magnetization, susceptibility, and correlation length. These equations yield consistent power-law behaviors near the critical point, with critical exponents expressed analytically in terms of conformable parameters. The formalism preserves dimensional consistency and remains compatible with equilibrium thermodynamics, while offering greater flexibility than standard approaches. In the limiting case $\mu\to 1$, the conformable derivative reduces to the standard derivative, and the classical Ginzburg-Landau theory with mean-field critical exponents is recovered.
In particular, the observed deviations from the mean-field exponents are explained as consequences of conformable deformation, capturing finite-size effects, spatial granularity, and critical slowing down. The flexibility of the deformation parameter enables the modeling of crossover behaviors between distinct universality classes, potentially capturing nontrivial fixed-point trajectories. The conformable approach thus provides a meaningful interpolation between classical mean-field theory and more general statistical frameworks, such as nonextensive thermodynamics.
We demonstrated the relevance of the conformal framework by fitting of experimental data of the superconducting transition of niobium. Good fits could be achieved over relatively broad intervals around the critical temperature for the heat capacity, the London penetration depth, and the coherence length. For this purpose, an asymmetric formulation was proposed based on the conformal GL-formulation, in which different model parameters appear below and above the critical temperature.
Contrary to the impression that the present formulation merely restates known mean-field scaling, our results highlight several nontrivial novel physical insights:
(i) Dynamical origin of static exponents. By combining the conformable derivative with the critical kinetic coefficient $\Gamma(T)\sim|T-T_{c}|^{z\nu}$, the static exponents emerge directly from the relaxation dynamics, rather than being postulated. For instance, the order parameter equation yields the form
$\Gamma(T)T^{1-\mu_{\psi}}\frac{d\psi}{dT}\simeq-a_{0}(T_{c}-T)\psi,$
leading to $\psi(T)\propto(T_{c}-T)^{\beta(\mu_{\psi})}$, where $\beta$ is determined by $(\mu_{\psi},\Gamma)$. This mechanism is absent in a simple reparametrizations of $T$.
(ii) Cross-observable constraints. Equation (104) establishes the internal consistency among the conformable parameters associated with different observables. Specifically, the exponents governing the temperature dependence of $\{C_{V},|M|,\chi,\xi\}$ are not independent but related by the classical scaling identities expressed in terms of the deformation indices $\mu_{X}$:
$\mu_{C}^{-1}=2\,\mu_{M}^{-1}+\mu_{\chi}^{-1}-2\,\mu_{\xi}^{-1}.$ (104)
Hence, once $\mu_{C}$ is determined experimentally, the remaining parameters $\{\mu_{M},\mu_{\chi},\mu_{\xi}\}$ are constrained by this relation. This feature defines a set of cross-observable tests for the conformable scaling hypothesis, providing a stringent criterion for internal consistency across thermodynamic observables.
(iii) Distinct experimental signatures. The prediction that $T^{\mu_{X}}d\ln X/dT$ has a constant slope in $\ln|T-T_{c}|$ provides a direct discriminant from conventional fits: the slope is observable-dependent and determined by microscopic couplings, not just by a universal change of variables.
(iv) Physical interpretation of $\mu$. The conformable index $\mu$ captures finite-size effects, disorder, and nontrivial memory of thermal fluctuations, bridging kinetic slowing down and effective fractal dimensionality. This extends beyond a notational reformulation.
A significant advantage of the conformable framework lies in its analytical tractability. Unlike renormalization group methods, which often rely on asymptotic expansions and perturbative schemes, the conformable model yields closed-form expressions that can be directly fitted to experimental data. This was demonstrated through application to superconducting phase transitions in niobium, where the model successfully captured the asymmetric scaling of specific heat and London penetration depth near the critical temperature. The extracted exponents are consistent with deviations from mean-field theory and suggest the presence of mesoscopic effects, spatial inhomogeneities, and memory-driven dynamics.
The geometric interpretation of the conformable framework developed herein is linked to deformations in the thermal or spatial metric Weberszpil and Helayël-Neto (2015); Rosa and Weberszpil (2018); Has et al. (2024) and offers insight into how fractal geometry and nonlocal responses might influence critical behavior. In parallel, conformable symmetry approaches have also found applications in nuclear structure and critical point symmetries Raissi-Rad et al. (2021), illustrating the broader relevance of such deformations across diverse physical systems. In particular, the framework is general and can be adapted to a broad range of critical systems, including percolation thresholds, spin glasses, disordered media, and quantum phase transitions. Its generality makes it a valuable tool for exploring universality in complex systems, particularly in non-equilibrium or nonextensive regimes.
Beyond phenomenological applications, several open questions remain as to the rigorous mathematical formulation of the conformable framework. These include the development of a well-defined operator semigroup structure, the spectral theory of deformed differential operators, and the formulation of variational principles compatible with conformable dynamics Godinho et al. (2020). Addressing these foundational issues may reveal deeper connections between the conformable approach and the broader structure of mathematical physics.
The possible extension of the conformal framework to quantum phase transitions, non-equilibrium critical dynamics, and systems with complex topology will be discussed in the future. These applications require further theoretical development and experimental validation.
Acknowledgements.
JW wishes to express their gratitude to FAPERJ, APQ1, for partial financial support. RM acknowledges the German Science Foundation (DFG, grant ME 1535/22-1) for support.
Declarations
Conflicts of interest/Competing interests:
The authors declare no conflicts of interest related to this work.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used ChatGpt in order to impriove the english. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
Appendix A Dimensional consistency
The mathematical structure of the conformable derivative suggests a deeper geometric meaning. In this section, we explore interpretations based on fractal metrics and spatial deformation. Specifically, we analyze the dimensional consistency of the primary equations and critical exponents derived using the conformable derivative framework.
A.1 Conformable derivative units
The conformable derivative with respect to temperature is defined as
$D_{T}^{(\mu)}f(T)=T^{1-\mu}\frac{df}{dT}$ (105)
Given that temperature has units of Kelvin, $T]=\text{K}$ we see that $\frac{df}{dT}$ has units of $f]/\text{K}$. Therefore
$D_{T}^{(\mu)}f]=[f]\cdot\mathrm{K}^{-\mu}.$ (106)
This confirms that the conformable derivative scales as expected, preserving consistent units when applied to temperature-dependent quantities.
A.2 Critical exponent dimensionality
The critical exponents $\alpha$, $\beta$, $\gamma$, and $\nu$ are defined through
$\alpha=\kappa T_{c}^{\mu_{C}-1},\quad\beta=\gamma T_{c}^{\mu_{M}-1},$
$\gamma=\lambda T_{c}^{\mu_{\chi}-1},\quad\nu=\rho T_{c}^{\mu_{\xi}-1}.$ (107)
Since exponents such as $\alpha$ appear in power-law expressions of the form $C(T)\sim|T-T_{c}|^{-\alpha}$, they must be dimensionless. For this to hold, $\kappa$ must have units of $\text{K}^{1-\mu_{C}}$, and similarly for the other exponents. Hence, with this choice all critical exponents are indeed dimensionless, satisfying the necessary conditions for physical consistency.
A.3 Modified Ginzburg–Landau equation
The modified GL-equation derived using the conformable kinetic term reads
$a(T)\psi+b|\psi|^{2}\psi-\frac{1}{2m}T^{2(1-\alpha)}\frac{d^{2}\psi}{dT^{2}}-\frac{1-\alpha}{m}T^{1-2\alpha}\frac{d\psi}{dT}=0.$ (108)
Now let $\psi$ have the units $\psi]=\text{K}^{-1/2}$. Then (i) $a(T)=a_{0}(T-T_{c})$ implies $a_{0}]=\text{K}^{-2}$; (ii) to match units, the coefficient $b$ must have the units $b]=\text{K}^{-5/2}$; (iii) the kinetic terms involve second and first derivatives, and with the conformable powers of $T$, they retain the same dimension as $\psi/\text{K}$, ensuring consistency. Thus, all terms in the equation have matching units.
A.4 Scaling of the heat capacity
The heat capacity scales as
$C(T)\sim|T-T_{c}|^{-\alpha},\quad[C]=\frac{E}{T}$ (109)
For dimensional consistency, $\alpha$ must be dimensionless, which is satisfied given that $\alpha=\kappa T_{c}^{\mu_{C}-1}$.
A.5 Penetration depth and order parameter
The London penetration depth scales as
$\lambda_{L}(T)\sim(T_{c}-T)^{-\zeta},\quad\zeta=\beta/2.$ (110)
Assuming that $\lambda_{L}]=\text{length}$, and that $\zeta$ is dimensionless, the units are correctly preserved.
In summary, all core expressions and critical exponent definitions derived from the conformable derivative framework are dimensionally consistent. This ensures the robustness of the theoretical framework and its applicability to thermodynamical systems such as superconducting transitions.
Appendix B Variational derivation of the modified Ginzburg-Landau equation
We start from the conformable free energy functional
$\mathcal{F}_{\alpha}[\psi]=\int\left[a(T)|\psi|^{2}+\frac{b}{2}|\psi|^{4}+\frac{1}{2m}\left|T_{\alpha}\psi(T)\right|^{2}\right]dT.$ (111)
The conformable derivative with respect to temperature reads
$T_{\alpha}\psi(T)=T^{1-\alpha}\frac{d\psi}{dT}.$ (112)
Then, the kinetic term takes on the form
$\left|T_{\alpha}\psi(T)\right|^{2}=T^{2(1-\alpha)}\left|\frac{d\psi}{dT}\right|^{2}$ (113)
and the full functional reads
$\mathcal{F}_{\alpha}[\psi]=\int\left[a(T)|\psi|^{2}+\frac{b}{2}|\psi|^{4}+\frac{1}{2m}T^{2(1-\alpha)}\left|\frac{d\psi}{dT}\right|^{2}\right]dT.$ (114)
To minimize the free energy, we compute the variation with respect to $\psi^{*}(T)$,
$\delta\mathcal{F}_{\alpha}$ $=$ $\int\left[a(T)\delta|\psi|^{2}+b|\psi|^{2}\delta|\psi|^{2}\right$. (115)
$\hskip-45.52458pt\left.+\frac{1}{2m}T^{2(1-\alpha)}\left(\frac{d\psi}{dT}\frac{d(\delta\psi^{*})}{dT}+\frac{d\psi^{*}}{dT}\frac{d(\delta\psi)}{dT}\right)\right]dT.$
Using integration by parts and assuming vanishing boundary terms,
$\int T^{2(1-\alpha)}\frac{d\psi}{dT}\frac{d(\delta\psi^{*})}{dT}dT$
$=-\int\delta\psi^{*}\left[\frac{d}{dT}\left(T^{2(1-\alpha)}\frac{d\psi}{dT}\right)\right]dT.$ (116)
Therefore, setting $\delta\mathcal{F}_{\alpha}=0$, the associated Euler–Lagrange equation becomes
$a(T)\psi+b|\psi|^{2}\psi-\frac{1}{2m}\frac{d}{dT}\left(T^{2(1-\alpha)}\frac{d\psi}{dT}\right)=0.$ (117)
Expanding the total derivative, we obtain
$a(T)\psi+b|\psi|^{2}\psi-\frac{1}{2m}T^{2(1-\alpha)}\frac{d^{2}\psi}{dT^{2}}$
$\hskip 22.76228pt-\frac{(1-\alpha)}{m}T^{1-2\alpha}\frac{d\psi}{dT}=0.$ (118)
Final form of in powers of $1-T/T_{c}$
Substituting the ansatz
$\psi(T)=\psi_{0}\left(1-\frac{T}{T_{c}}\right)^{\beta}$ (119)
into the full equation, we obtain
$a_{0}(T-T_{c})\psi_{0}\left(1-\frac{T}{T_{c}}\right)^{\beta}+b\psi_{0}^{3}\left(1-\frac{T}{T_{c}}\right)^{3\beta}$
$\quad-\frac{1}{2m}T^{2(1-\alpha)}\frac{\beta(\beta-1)\psi_{0}}{T_{c}^{2}}\left(1-\frac{T}{T_{c}}\right)^{\beta-2}$
$\quad+\frac{(1-\alpha)}{m}T^{1-2\alpha}\frac{\beta\psi_{0}}{T_{c}}\left(1-\frac{T}{T_{c}}\right)^{\beta-1}=0.$ (120)
Matching powers of $(1-T/T_{c})$, we recover $\beta=1/2$ as the critical exponent.
Now, we analyze the behavior of $\psi(T)$ near the critical temperature $T_{c}$. To this end, consider again the nonlinear temperature-deformed GL-type equation
$a(T)\psi+b|\psi|^{2}\psi-\frac{1}{2m}T^{2(1-\alpha)}\frac{d^{2}\psi}{dT^{2}}$
$-\frac{1-\alpha}{m}T^{1-2\alpha}\frac{d\psi}{dT}=0,$ (121)
with $a(T)=a_{0}(T-T_{c})$. We again assume the ansatz
$\psi(T)=\psi_{0}\left(1-\frac{T}{T_{c}}\right)^{\beta}=\psi_{0}\varepsilon^{\beta},\quad\text{where }\varepsilon=1-\frac{T}{T_{c}}.$ (122)
Then, forming the derivatives with respect to T, we have
$\frac{d\varepsilon}{dT}=-\frac{1}{T_{c}},$
$\frac{d\psi}{dT}=-\frac{\beta\psi_{0}}{T_{c}}\varepsilon^{\beta-1}$
$\frac{d^{2}\psi}{dT^{2}}=\frac{\beta(\beta-1)\psi_{0}}{T_{c}^{2}}\varepsilon^{\beta-2}.$ (123)
We then obtain the different terms. Starting with the linear term,
$a_{0}(T-T_{c})\psi=-a_{0}T_{c}\psi_{0}\varepsilon^{\beta+1};$ (124)
then the nonlinear term
$b|\psi|^{2}\psi=b\psi_{0}^{3}\varepsilon^{3\beta};$ (125)
the term with the second derivative,
$\frac{1}{2m}T^{2(1-\alpha)}\frac{\beta(\beta-1)\psi_{0}}{T_{c}^{2}}\varepsilon^{\beta-2};$ (126)
and the term with the first-order derivative,
$\frac{(1-\alpha)\beta\psi_{0}}{mT_{c}}T^{1-2\alpha}\varepsilon^{\beta-1}.$ (127)
We see that $\varepsilon=1-T/T_{c}$ enters these different terms with different exponent. Seeking a consistent balance, we suppose that the nonlinear term $\propto\varepsilon^{3\beta}$ and the linear term $\propto\varepsilon^{\beta+1}$ dominate. We then equate their powers: $3\beta=\beta+1$, implying $\beta=1/2$. This is consistent with standard Landau theory and a typical critical exponent in second-order phase transitions. This means that in the four terms above, we have the orders
$\begin{aligned} &\varepsilon^{\beta-2}=\varepsilon^{-3/2}\\ &\varepsilon^{\beta-1}=\varepsilon^{-1/2}\\ &\varepsilon^{3\beta}=\varepsilon^{3/2}\\ &\varepsilon^{\beta+1}=\varepsilon^{3/2}\end{aligned},$ (128)
i.e., the derivative terms diverge as $\varepsilon\to 0$. However, despite their divergence, the derivative terms describe kinetic effects (gradients of $\psi$). Near $T=T_{c}$, the order parameter becomes small and smooth, $\psi\to 0$, and its variation is slow. Finally the mean-field and equilibrium theories focus on minimizing the free energy—gradient terms are subleading. This allows us to match
$a(T)\psi+b\psi^{3}=0\quad\Rightarrow\quad\psi=\pm\sqrt{-\frac{a(T)}{b}}.$ (129)
From this we conclude that
$\psi(T)=\psi_{0}\left(1-\frac{T}{T_{c}}\right)^{1/2}$ (130)
Thus, the derivative terms represent corrections to the scaling behavior, but not to the leading-order critical exponent. This demonstrates that the ansatz (119) satisfies the nonlinear equation near $T=T_{c}$, with leading order balance provided by the algebraic terms,
$a(T)\psi+b\psi^{3}=0\quad\Rightarrow\quad\boxed{\beta=1/2}.$ (131)
This matches classical Landau theory and remains valid under conformable kinetic corrections. Moreover, this represents a static equilibrium solution for which the derivatives are zero (i.e., long-time equilibrium). It is important to note that our model for equilibrium and homogeneous conditions and when fluctuations are negligible, has the classical exponent $\beta=1/2$.
Appendix C Recovery of the relaxation equation in the undeformed limit
We now show that the classical critical exponent $\beta=\tfrac{1}{2}$ naturally emerges from the conformable relaxation equation in the undeformed equilibrium limit. Consider the generalized relaxation model
$\Gamma(T)\,T^{1-\mu_{\psi}}\frac{d\psi}{dT}=-a_{0}(T_{c}-T)\psi-b|\psi|^{2}\psi,$ (132)
where $\Gamma(T)$ is a temperature-dependent kinetic coefficient, $\mu_{\psi}\in(0,1]$ is the conformable deformation parameter, and the right-hand side arises from the variational derivative of the standard GL-free energy.
In the undeformed limit, we take
$\mu_{\psi}=1,\quad\text{so that}\quad T^{1-\mu_{\psi}}=1,$ (133)
and for simplicity we assume that the kinetic coefficient is constant near $T_{c}$, i.e., $\Gamma(T)\approx\Gamma_{0}$. Equation (132) then becomes
$\Gamma_{0}\frac{d\psi}{dT}=-a_{0}(T_{c}-T)\psi-b|\psi|^{2}\psi.$ (134)
We now look for a solution near the critical point of the form
$\psi(T)=\psi_{0}(T_{c}-T)^{\beta},$ (135)
valid for $T<T_{c}$ and where $\psi_{0}$ is a constant amplitude to be determined. Computing the derivative, we have
$\frac{d\psi}{dT}=-\beta\psi_{0}(T_{c}-T)^{\beta-1}.$ (136)
Substituting into Eq. (134), we obtain
$-\Gamma_{0}\beta\psi_{0}(T_{c}-T)^{\beta-1}$
$=-a_{0}(T_{c}-T)\psi_{0}(T_{c}-T)^{\beta}-b\psi_{0}^{3}(T_{c}-T)^{3\beta}.$ (137)
Divide both sides by $\psi_{0}(T_{c}-T)^{\beta-1}$ (assuming $\psi_{0}\neq 0$ and $T<T_{c}$),
$\Gamma_{0}\beta=-a_{0}(T_{c}-T)^{2}-b\psi_{0}^{2}(T_{c}-T)^{2\beta+1}.$ (138)
Now we check the leading-order behavior as $T\to T_{c}^{-}$: The term $(T_{c}-T)^{2}$ dominates, while the nonlinear term $\propto(T_{c}-T)^{2\beta+1}$ is subleading if $2\beta+1>2$ and thus $\beta>\tfrac{1}{2}$; or they are comparable if $\beta=\tfrac{1}{2}$.
Assuming that the dominant balance occurs between the terms of order
$\Gamma_{0}\beta\approx-a_{0}(T_{c}-T)^{2},$ (139)
we see that this leads to a contradiction: the left-hand side is constant, while the right-hand side vanishes as $T\to T_{c}$. Therefore, this balance fails. Instead, if we assume that the dominant balance occurs between the first-derivative term and the nonlinear cubic term,
$\Gamma_{0}\beta\psi_{0}(T_{c}-T)^{\beta-1}\approx-b\psi_{0}^{3}(T_{c}-T)^{3\beta},$ (140)
and we cancel signs and factors of $\psi_{0}$, we see that
$\Gamma_{0}\beta(T_{c}-T)^{\beta-1}=b\psi_{0}^{2}(T_{c}-T)^{3\beta}.$ (141)
Divide both sides by $(T_{c}-T)^{\beta-1}$ to find
$\Gamma_{0}\beta=b\psi_{0}^{2}(T_{c}-T)^{2\beta+1}.$ (142)
To keep both sides finite and nonzero as $T\to T_{c}$, we must have
$2\beta+1=0\quad\Rightarrow\quad\beta=-\tfrac{1}{2},$ (143)
which is unphysical.
So we try instead to match the linear and nonlinear terms (equivalent to the adiabatic limit),
$a_{0}(T_{c}-T)\psi\sim b\psi^{3}\quad\Rightarrow\quad a_{0}(T_{c}-T)\sim b\psi^{2}.$ (144)
Solving for $\psi(T)$, we see that
$\psi^{2}(T)\sim\frac{a_{0}}{b}(T_{c}-T),\quad\Rightarrow\quad\psi(T)\sim\sqrt{\frac{a_{0}}{b}}(T_{c}-T)^{1/2}.$ (145)
Thus, the leading-order balance occurs not with the derivative term, but with the free-energy terms, confirming that in the long-time (quasi-equilibrium) limit, the stationary solution satisfies
$\beta=\frac{1}{2}.$ (146)
This demonstrates that even when starting from the generalized relaxation equation, the classical mean-field exponent $\beta=1/2$ is recovered in the undeformed limit $\mu_{\psi}=1$, under equilibrium assumptions. The derivative terms decay near $T_{c}$, and the critical scaling is governed by the balance of the potential terms in the GL-free energy.
Higher-order corrections
Derivative terms can be treated as perturbations (next-to-leading order). They modify the scaling only slightly and do not affect the critical exponent to leading order. Thus, (i) $\varepsilon^{\beta-1}$ and $\varepsilon^{\beta-2}$ diverge as $\varepsilon\to 0$ but are derivative-based; (ii) near $T_{c}$, $\psi$ is small and smooth, so derivatives are small; (iii) the dominant balance is between $a(T)\psi$ and $b\psi^{3}$, yielding again $\beta=1/2$. Thus, indeed the derivative terms are corrections to this leading-order behavior.
Appendix D On the claim of simple reparametrization
For a scalar $f(T)$ with constant coefficients, we have that $D_{T}^{(\mu)}f=T^{1-\mu}f^{\prime}(T)=df/dT^{\prime}$ with $T^{\prime}=T^{\mu}/\mu$. However, consider the two cases:
(i) Critical kinetics. With $\Gamma(T)\!\sim\!|T-T_{c}|^{z\nu}$, the relaxation law becomes $\Gamma(T(T^{\prime}))\,\tfrac{dX}{dT^{\prime}}=-\partial F/\partial X$. The nontrivial $T$-dependence of $\Gamma$ survives, producing a distinct scaling.
(ii) Thermal GL-functional. The conformable GL-free energy includes a gradient term $\propto T^{2(1-\alpha)}|\tfrac{d\psi}{dT}|^{2}$. Under a transformation $T\mapsto T^{\prime}$, Jacobian factors remain, yielding Euler-Lagrange equations [Eq. (118)] that are not reducible to the standard GL-form.
Hence, the conformable framework only collapses to a trivial change of variables in the static, constant-coefficient case. In the presence of critical kinetics and weighted GL-terms, it carries genuine physical content.